### v1

# Extremal random matrices with independent entries and matrix superconcentration inequalities

Abstract.
We prove nonasymptotic matrix concentration inequalities for the spectral norm of (sub)gaussian random matrices with centered independent entries that capture fluctuations at the Tracy-Widom scale. This considerably improves previous bounds in this setting due to Bandeira and Van Handel, and establishes the best possible tail behavior for random matrices with an arbitrary variance pattern. These bounds arise from an extremum problem for nonhomogeneous random matrices: among all variance patterns with a given sparsity parameter, the moments of the random matrix are maximized by block-diagonal matrices with i.i.d. entries in each block. As part of the proof, we obtain sharp bounds on large moments of Gaussian Wishart matrices.
Key words and phrases:
Random matrices, spectral norm, matrix concentration inequalities
2010 Mathematics Subject Classification:
60B20; 60E15; 46L53; 46L54; 15B52
1. Introduction
Let $X$ be an $n\times m$ matrix whose entries are independent, centered Gaussian variables with an arbitrary variance pattern $X_{ij}\sim N(0,b_{ij}^{2})$. This paper is concerned with bounding the spectral norm $\|X\|$ of such matrices.
In the special case that $b_{ij}=1$ for all $i,j$, we recover one of the most classical models of random matrix theory: then $\|X\|^{2}=\|X^{*}X\|$ is the norm of a Wishart matrix with unit covariance. In this case, classical methods of random matrix theory yield the exact asymptotic behavior [11]
$\lim_{\begin{subarray}{c}n,m\to\infty\\ m=cn\end{subarray}}\mathbf{P}\big{[}\|X\|>\sqrt{n}+\sqrt{m}+\tfrac{1}{2}(\tfrac{1}{\sqrt{n}}+\tfrac{1}{\sqrt{m}})^{1/3}s\big{]}=1-F_{1}(s),$ (1.1)
where $F_{1}(s)$ is the distribution function of the Tracy-Widom law of order one. From this expression, we immediately read off that in the above asymptotic regime $\|X\|=(1+o(1))(\sqrt{n}+\sqrt{m})$ with fluctuations of order $n^{-1/6}\vee m^{-1/6}$. It is important to note that the scale of the fluctuations is much smaller than what is predicted by general concentration of measure principles [13], which yield fluctuations of order $O(1)$ in the present setting. The presence of such unexpectedly small fluctuations is sometimes called the superconcentration phenomenon [6].
The Wishart model is amenable to explicit computations due to its simple structure and large degree of symmetry. However, there is little hope to perform explicit computations for an arbitrary variance pattern $(b_{ij})$, as such models can exhibit a wide variety of different structures and behaviors that are specific to the given pattern (for example, this class includes random band matrices and sparse matrices with an arbitrary deterministic sparsity pattern as special cases). Nonetheless, given their importance in many applications, it is of considerable interest to obtain bounds for random matrices with an arbitrary variance pattern.
One of the main results in this direction was obtained some years ago by Bandeira and the second author. To describe this result, define the parameters
$\sigma_{1}^{2}:=\max_{j\leq m}\sum_{i\leq n}b_{ij}^{2},\qquad\quad\sigma_{2}^{2}:=\max_{i\leq n}\sum_{j\leq m}b_{ij}^{2},\qquad\quad\sigma_{*}^{2}:=\max_{\begin{subarray}{c}i\leq n\\ j\leq m\end{subarray}}b_{ij}^{2}.$ (1.2)
The following is shown in [3, Corollary 3.11].
Theorem 1.1 ([3]).
Let $n\leq m$, and let $X$ be the $n\times m$ random matrix whose entries $X_{ij}\sim N(0,b_{ij}^{2})$ are independent. Then we have
$\mathbf{P}\big{[}\|X\|>(1+\varepsilon)(\sigma_{1}+\sigma_{2})+\sigma_{*}t\big{]}\leq n\,e^{-C_{\varepsilon}t^{2}}$
for all $t\geq 0$ and $0<\varepsilon\leq\frac{1}{2}$, where the constant $C_{\varepsilon}$ depends only on $\varepsilon$.
Theorem 1.1 implies that
[…]
In the Wishart case $b_{ij}=1$ for all $i,j$, this yields $\|X\|\leq(1+o(1))(\sqrt{n}+\sqrt{m})$ which captures the exact leading order behavior in (1.1). However, the order of the fluctuations in Theorem 1.1 is much larger than in (1.1): the second term in (1.3) diverges as $n,m\to\infty$, while the second order term in (1.1) is $o(1)$. In particular, Theorem 1.1 fails to recover any form of superconcentration.
On the other hand, the large second term in (1.3) is unavoidable in general for sparse matrices, as the following classical example illustrates.
Example 1.2.
Let $n=m$ and $b_{ij}=1_{|i-j|\leq k}$, that is, $X$ is a random band matrix with bandwidth $12k+$ 1. Then we have
$\sigma_{1}+\sigma_{2}=2\sqrt{2k+1},\qquad\quad\|X\|\geq\max_{i}|X_{ii}|=(1+o(1))\sqrt{2\log n}$
with high probability as $n\to\infty$. It follows that the second term in (1.3) dominates the behavior of $\|X\|$ in the sparse regime $k\ll\log n$. In fact, (1.3) optimally captures the phase transition for the leading order behavior of the norm of random band matrices (cf. [3, Corollary 4.4] in the self-adjoint case).
Such examples suggest that the large scale of the fluctuations in Theorem 1.1 is a necessary feature of any bound for random matrices with an arbitrary variance pattern. Surprisingly, this expectation turns out to be incorrect. The main results of this paper will yield a considerable improvement on Theorem 1.1, which simultaneously captures the fluctuations at Tracy-Widom scale of (1.1) and sharpens the phase transition between sparse and dense matrices that is implicit in (1.3).
1.1. Main results
While the main results of this paper will be proved both for non-self-adjoint and for self-adjoint random matrices, we focus the presentation in the introduction on the non-self-adjoint case for concreteness. We further consider a slightly more general setting than the above Gaussian model.
Model 1.3.
$X$ is an $n\times m$ matrix with $X_{ij}=b_{ij}\xi_{ij}$, where $b_{ij}\geq 0$ are arbitrary scalars and $\xi_{ij}$ are independent symmetrically distributed real random variables with $\mathbf{E}[\xi_{ij}^{2p}]\leq\mathbf{E}[g^{2p}]$ for all $i,j$ and $p\in\mathbb{N}$ (here $g\sim N(0,1)$).
In the setting of Model 1.3, we always define $\sigma_{1},\sigma_{2},\sigma_{*}$ as in (1.2).
1.1.1. Small deviations
Our main result in this setting is the following.
Theorem 1.4 (Small deviations).
Let $X$ be as in Model 1.3 and $\sigma_{1}\leq\sigma_{2}$. Then
$\mathbf{P}\big{[}\|X\|>\sigma_{1}+\sigma_{2}+\sigma_{*}^{4/3}\sigma_{1}^{-1/3}t\big{]}\leq\frac{n\sigma_{*}^{2}}{C\sigma_{1}^{2}}\,e^{-Ct^{3/2}}$
for all $0\leq t\leq\frac{\sigma_{1}^{1/3}\sigma_{2}}{\sigma_{*}^{4/3}}$, where $C$ is a universal constant.
Remark 1.5.
The assumption $\sigma_{1}\leq\sigma_{2}$ entails no loss of generality, as in the opposite case $\sigma_{1}>\sigma_{2}$ we may simply apply Theorem 1.4 to the adjoint matrix $X^{*}$.
In the Wishart case $b_{ij}=1$ for all $i,j$ with $n\leq m$, Theorem 1.4 yields
$\mathbf{P}\big{[}\|X\|>\sqrt{n}+\sqrt{m}+n^{-1/6}t\big{]}\leq C^{-1}e^{-Ct^{3/2}}$
for $0\leq t\leq n^{1/6}\sqrt{m}$. This captures precisely the fluctuations and the upper tail of the Tracy-Widom asymptotics (1.1), up to a universal constant. To the best of our knowledge, the nonasymptotic bound is new even in this very special case.
For arbitrary variance patterns with $\sigma_{1}\leq\sigma_{2}$, Theorem 1.4 yields
[…]
provided $\sigma_{1}^{1/4}\sigma_{2}^{3/4}\gtrsim\sigma_{*}\sqrt{\log n}$. In particular, as soon as $\sigma_{1}\gg\sigma_{*}\sqrt{\log n}$, (1.4) improves drastically on (1.3). We will explain in section 1.1.3 that this bound is essentially the best possible even for nonhomogeneous random matrices.
1.1.2. Large deviations
Theorem 1.4 controls small deviations, that is, deviations up to the order of the mean. In contrast, the large deviations of $\|X\|$ are controlled by its Gaussian tail behavior [17, Corollary 3.2], and we cannot expect a qualitative improvement over Theorem 1.1 in this setting.
Nonetheless, the basic principle behind Theorem 1.4 gives rise to a quantitative improvement: a variant of Theorem 1.1 with optimal constants.
Theorem 1.6 (Large deviations).
Let $X$ be as in Model 1.3 and $n\leq m$. Then
$\mathbf{P}\big{[}\|X\|>\sigma_{1}+\sigma_{2}+\sigma_{*}(1+t)\big{]}\leq 2n\,e^{-t^{2}/2}$
for all $t\geq 0$.
Theorem 1.6 implies that
[…]
as $n\to\infty$. The significance of this result is that in many examples, the second term in (1.5) matches the trivial lower bound $\|X\|\geq\max_{i,j}|X_{ij}|$; this is the case, for example, for the random band matrix of Example 1.2. In such situations, (1.5) captures the exact leading order behavior $\|X\|=(1+o(1))\sigma_{*}\sqrt{2\log n}$ in the sparse regime $\sigma_{1}+\sigma_{2}\ll\sigma_{*}\sqrt{\log n}$, while (1.3) necessarily loses a universal constant.
1.1.3. Extremum principle
The proofs of Theorems 1.4 and 1.6 are based on a more fundamental principle that is of independent interest.
To explain this principle, let us begin by recalling the idea behind the proof of Theorem 1.1. Rather than bound the norm of the nonhomogeneous model $X$ directly, the approach of [3] compares $X$ with an i.i.d. matrix of smaller dimension, whose norm can be controlled by any classical method for homogeneous random matrices. In particular, the main technical device in [3] shows that when normalized so that $\sigma_{*}=1$, the $2p$-moment of $X$ can be bounded by the $2p$-moment of a matrix with i.i.d. entries of dimension $\lceil\sigma_{1}^{2}+p\rceil\times\lceil\sigma_{2}^{2}+p\rceil$. This suffices to capture the leading order behavior of $\|X\|$, but the dependence of the comparison matrix on $p$ precludes any accurate control of the fluctuations.
The basis for the results of this paper may be viewed as an optimal comparison principle of this kind. In the following, we denote by $\mathop{\mathrm{Tr}}M$ the (unnormalized) trace and by $\mathop{\mathrm{tr}}M:=\frac{1}{n}\mathop{\mathrm{Tr}}M$ the normalized trace of an $n\times n$ matrix $M$.
Theorem 1.7 (Extremum principle).
Let $X$ be as in Model 1.3 with $\sigma_{*}=1$. Then
$\mathbf{E}[\mathop{\mathrm{tr}}(XX^{*})^{p}]\leq\mathbf{E}[\mathop{\mathrm{tr}}(YY^{*})^{p}]\quad\text{for all }p\in\mathbb{N},$
where $Y$ is the $\lceil\sigma_{1}^{2}\rceil\times\lceil\sigma_{2}^{2}\rceil$ matrix with independent entries $Y_{ij}\sim N(0,1)$.
Note that the inequality in Theorem 1.7 holds with equality when $X$ has i.i.d. entries. Thus, in contrast to the comparison principle of [3], Theorem 1.7 may be viewed as an extremum principle for random matrices. In terms of normalized moments, this principle may be expressed as follows.
Corollary 1.8.
Fix $d_{1},d_{2}\in\mathbb{N}$. Among all $X$ as in Model 1.3 with $\sigma_{*}^{2}=1$, $\sigma_{1}^{2}\leq d_{1}$, $\sigma_{2}^{2}\leq d_{2}$, and arbitrary dimensions $n\times m$, the normalized moments $\mathbf{E}[\mathop{\mathrm{tr}}(XX^{*})^{p}]$ are maximized for all $p$ by the matrix of dimension $d_{1}\times d_{2}$ with i.i.d. $N(0,1)$ entries.
In terms of unnormalized moments, we obtain the following.
Corollary 1.9.
Fix $n,m,d_{1},d_{2}\in\mathbb{N}$ with $\frac{m}{d_{2}}\geq\frac{n}{d_{1}}\in\mathbb{N}$. Among all $X$ as in Model 1.3 with $\sigma_{*}^{2}=1$, $\sigma_{1}^{2}\leq d_{1}$, $\sigma_{2}^{2}\leq d_{2}$, and fixed dimensions $n\times m$, the unnormalized moments $\mathbf{E}[\mathop{\mathrm{Tr}}(XX^{*})^{p}]$ are maximized for all $p$ by the block-diagonal matrix whose blocks have dimension $d_{1}\times d_{2}$ and i.i.d. $N(0,1)$ entries (Figure 1.1).
$\scriptstyle n$$\scriptstyle m$$\scriptstyle d_{1}$$\scriptstyle d_{2}$
Figure 1.1. An extremal block-diagonal matrix.
In particular, that block-diagonal matrices have the largest moments explains the form of Theorem 1.4. Indeed, as the norm of a block-diagonal matrix is the maximum of the norms of its blocks, (1.1) suggests that its distribution is approximately the maximum of $\frac{n}{d_{1}}$ independent Tracy-Widom variables. The tail probabilities of this distribution are precisely of the form that is captured by the tail bound of Theorem 1.4. In particular, the bound (1.4) is essentially the best possible in that it yields the correct behavior of block-diagonal matrices.
Theorem 1.7 will be proved in section 4 below. With this result in hand, the proof of Theorems 1.4 and 1.6 will reduce to the proof of analogous tail bounds for Wishart matrices. The latter is made possible by a beautiful method, pioneered by Ledoux [14, 15, 16], for deriving Tracy-Widom type tail bounds from sharp moment estimates. While such estimates were obtained by Ledoux for GUE [14] and GOE [16] matrices, sharp moment bounds for Wishart matrices do not appear to be known in the literature. Such bounds are obtained here in section 5. The Wishart model is considerably more delicate than the GUE and GOE models due to the fact that we require bounds that hold uniformly for all dimensions $n,m$, not just in the asymptotic regime (1.1) where $m=cn$ are proportional.
Let us emphasize that, despite the apparent similarity between Theorem 1.7 and the comparison principles of [3], the proofs of these results are completely different. The method of [3] uses a classical combinatorial expression for the moments as a sum over equivalence classes of even closed walks, and estimates each term separately. In contrast, the proof of Theorem 1.7 expresses the moments of a Gaussian random matrix as a sum over pairings by the Wick formula, and then uses an iterative procedure to reduce each term to a noncrossing pairing. This new approach turns out to provide a much more efficient mechanism for controlling the moments.
1.1.4. Self-adjoint models and outline
While we have focused the discussion in the introduction on non-self-adjoint models for concreteness, analogous results hold also for self-adjoint models. Beside that self-adjoint random matrices arise frequently in applications, the proofs of the self-adjoint analogues of our main results turn out to be simpler than those in the non-self-adjoint case. We will therefore develop these results in detail before we return to the non-self-adjoint case.
  1. $\bullet$
    In section 2 we consider self-adjoint random matrices with independent complex Gaussian entries. The extremum principle admits a particularly simple proof in this setting that avoids almost all the complications that arise in the real case.
  2. $\bullet$
    We then consider in section 3 self-adjoint random matrices with independent real entries, that is, the self-adjoint analogue of Model 1.3. The proof of the the extremum principle is much more involved in this case, and shares the same difficulties as the proof of the extremum principle for non-self-adjoint models. However, for both complex and real self-adjoint models, we can directly invoke the moment estimates of Ledoux [14, 15, 16] to obtain tail bounds.
  3. $\bullet$
    Finally, the non-self-adjoint case is developed in section 4, while section 5 developes the sharp moment estimates that are needed in this case.
1.2. Discussion and open questions
1.2.1. Intrinsic freeness
To date, two main approaches have been developed for obtaining sharp nonasymptotic bounds for the spectral norm of nonhomogeneous random matrices $X$ that are of a fundamentally different nature:
  1. 1.
    The extremum principles of [3] and of this paper compare the spectral statistics of $X$ with those of an i.i.d. random matrix $Y$ of smaller dimension.
  2. 2.
    The intrinsic freeness principle of [1] compares the spectral statistics of $X$ with those of a deterministic operator $X_{\rm free}$ that arises in free probability theory.
The theory of [1] is in fact much more general than the setting of this paper, in that it captures both random matrices with dependent entries and more general spectral statistics. However, the following special case may be directly compared to the bounds of this paper: if $X$ is an $n\times m$ random matrix with independent Gaussian entries $X_{ij}\sim N(0,b_{ij}^{2})$ and $n\leq m$, then [1, Corollary 2.2] yields
[…]
for every $\varepsilon>0$ (this may be shown as in [1, Lemma 3.1]). As
$\|X_{\rm free}\|\leq\sigma_{1}+\sigma_{2}$
by [1, Lemma 2.5], the intrinsic freeness principle readily yields a slightly weaker form of (1.3). Note that the second term in (1.6) is even larger than that of (1.3), while the results of this paper yield a much smaller second-order term.
On the other hand, there are many situations where the intrinsic freeness principle yields strictly better results than those of this paper: whenever
$\sigma_{*}(\log m)^{3/2}\ll\|X_{\rm free}\|\leq(1-\delta)(\sigma_{1}+\sigma_{2})$
for some $\delta>0$, the bound of (1.6) is strictly better to leading order, which renders any improvement to the second-order term negligible.
In view of the above discussion, the bounds of this paper are of particular interest when $\|X_{\rm free}\|=\sigma_{1}+\sigma_{2}$. The following lemma explains when this is the case.
Lemma 1.10.
$\|X_{\rm free}\|=\sigma_{1}+\sigma_{2}$ when $\sum_{i}b_{ij}^{2}=\sigma_{1}^{2}$ and $\sum_{j}b_{ij}^{2}=\sigma_{2}^{2}$ for all $i,j$.
Proof.
We can apply [1, Lemma 3.2 and Remark 2.6] to obtain
$\|X_{\rm free}\|\geq 2\sum_{i}\sqrt{w_{i}\sum_{j}b_{ij}^{2}v_{j}}+2\sum_{j}\sqrt{w_{i}\sum_{i}b_{ij}^{2}v_{j}}$
for every $w\in\mathbb{R}^{n}_{+}$, $v\in\mathbb{R}^{m}_{+}$ with $\sum_{i}w_{i}+\sum_{j}v_{j}=1$. The conclusion follows by choosing $w_{i}=\frac{1}{2n}$, $v_{j}=\frac{1}{2m}$, and noting that $\sum_{i,j}b_{ij}^{2}=m\sigma_{1}^{2}=n\sigma_{2}^{2}$. ∎
Variance patterns with constant row and column sums as in Lemma 1.10 arise naturally in applications, for example, in the study of sparse random matrices whose sparsity pattern is biregular. At the same time, the results of this paper are stronger than those obtained by the intrinsic freeness principle for very sparse matrices (so that they capture phase transitions as in Example 1.2), and provide the strongest easily computable norm bounds for arbitrary variance patterns.
One may wonder whether it might in fact be possible to achieve the best of both worlds: could the tail bound of Theorem 1.4 with fluctuations at Tracy-Widom scale remain valid if the leading term $\sigma_{1}+\sigma_{2}$ is replaced by $\|X_{\rm free}\|$? The following example shows that such a bound cannot exist.
Example 1.11.
Let $n=m$ and $b_{ij}^{2}=n^{-1}(1+\delta^{2}1_{i=1})$, known as the spiked covariance model (originally due to Baik, Ben Arous, and Péché in the complex case). Then the leading order behavior of $\|X\|$ exhibits the following phase transition:
[…]
as $n\to\infty$ (cf. [1, 2] and [19, §2.1]). Moreover, it is shown in [19, Theorem 3] that when $\delta>1$, the fluctuations of $\|X\|$ are of order $\sigma_{*}\sim n^{-1/2}$. This shows that we cannot replace $\sigma_{1}+\sigma_{2}$ by $\|X_{\rm free}\|$ in Theorem 1.4, as that would imply a much smaller bound on the fluctuations of order $\sigma_{*}^{4/3}$.
Example 1.11 illustrates in a particularly clear manner that the extremum and intrinsic freeness principles capture fundamentally different mathematical phenomena. This may appear somewhat surprising, given that both the proofs in this paper and those of [1] are based on the control of crossings in the Wick formula (in the latter case, this idea is due to [22]). However, these results exploit crossings in very different ways: here we show that the contribution of each crossing to the nonhomogeneous model is bounded by that in the i.i.d. model, while the intrinsic freeness principle aims to show that crossings are negligible altogether.
1.2.2. Superconcentration
The concentration of measure phenomenon [13] is a powerful tool for bounding the fluctuations of random structures. While general concentration inequalities often yield optimal bounds, there are also situations where the scale of the flucutations is much smaller than is predicted by general principles. This phenomenon was called superconcentration in [6].
When applied to the random matrix models of this paper, classical concentration inequalities ensure that the scale of the fluctuations of $\|X\|$ is at most of order $\sigma_{*}$, cf. [1, Corollary 4.14]. In contrast, the fluctuation term in Theorem 1.4, of order $\sigma_{*}^{4/3}\sigma_{1}^{-1/3}$, is often much smaller than $\sigma_{*}$. We may therefore think of Theorem 1.4 as a “matrix superconcentration inequality”. Indeed, this bound accurately captures the well known superconcentration property of the upper tail of the norm of Wishart matrices, and extends it to a much larger class of models. [^1] After the completion of this paper, we learned that a weaker superconcentration property was previously obtained by entirely different methods in [9, Theorem 3.4] for self-adjoint random matrices satisfying the analogue of Lemma 1.10. Theorem 3.4 below yields a much more precise tail bound in the self-adjoint setting that is essentially optimal by the extremum principle.
This interpretation must be treated with care, however, as the second-order term in Theorem 1.4 can only provide information on the fluctuations of $\|X\|$ when the first-order term is captured correctly. In particular, Theorem 1.4 can only establish genuine superconcentration for variance patterns that satisfy conditions as in Lemma 1.10. This is necessarily the case: Example 1.11 illustrates that general variance patterns may not exhibit any superconcentration at all.
More generally, it is not expected that the specific setting of Lemma 1.10 is necessary for superconcentration. General principles explained in [6] suggest that the spectral norm should exhibit superconcentration provided there are many singular values near the maximal one (this fails in Example 1.11, where the largest singular value is isolated from the bulk). Furthermore, even in the setting of Lemma 1.10, our results yield an upper bound that is sharp for block-diagonal matrices, but other such models may exhibit even smaller fluctuations (e.g., [21]). A precise understanding of when and how much superconcentration arises for nonhomogeneous random matrices remains out of reach of any known method.
1.2.3. Universality
The setting of Model 1.3 requires the entries of $X$ to be symmetrically distributed and have all their moments dominated by those of the Gaussian distribution. However, the classical Tracy-Widom asymptotics (1.1) remain valid in a much more general setting [20]: it suffices that the entries of $X$ have zero mean, unit variance, and soft control of the higher moments.
It is of considerable interest to understand whether nonasymptotic bounds can also be achieved in this much more general setting. An extension of Theorem 1.1 along these lines was proved in [12, §4.3]: it was shown there that for any random matrix $X$ whose entries are independent and have zero mean, the statement of Theorem 1.1 remains valid if we replace the parameters (1.2) by
$\sigma_{1}^{2}\leftarrow\max_{j}\sum_{i}\mathrm{Var}(X_{ij}),\qquad\sigma_{2}^{2}\leftarrow\max_{i}\sum_{j}\mathrm{Var}(X_{ij}),\qquad\sigma_{*}^{2}\leftarrow\max_{i,j}\|X_{ij}\|_{\infty}^{2}.$
However, just as Theorem 1.1, this result fails to capture fluctuations at the Tracy-Widom scale. It is natural to conjecture the validity of an analogous extension of Theorem 1.4 and of the other main results of this paper. Such results would be of particular interest in applications, e.g., to random graphs [4].
Unfortunately, it is not clear whether the methods of this paper can be adapted to achieve such results (nor do they follow from the very general universality principles in [5], whose error terms are far larger than the fluctuations in Theorem 1.4). While much less precise, the method used in [3, 12] does not use any special properties of the Gaussian distribution and is therefore readily adapted to more general models. In contrast, the proof of the sharp extremum principle of Theorem 1.7 is based on the Wick formula, and is therefore inherently Gaussian in nature. For this reason, an analogous extension of our main results remains an open problem.
1.3. Notation
The following notation will be used throughout this paper.
We denote by $N(0,\sigma^{2})$ or by $N_{\mathbb{R}}(0,\sigma^{2})$ the distribution of a real Gaussian random variable with mean 00 and variance $\sigma^{2}$. The distribution of a complex random variable, whose real and imaginary parts are independent with distribution $N_{\mathbb{R}}(0,\frac{\sigma^{2}}{2})$ (i.e., a complex Gaussian variable), is denoted as $N_{\mathbb{C}}(0,\sigma^{2})$.
For a matrix $M$, its adjoint is denoted as $M^{*}$; in particular, $a^{*}$ denotes the complex conjugate of $a\in\mathbb{C}$. We always denote by $\|M\|$ the spectral norm (i.e., the largest singular value) of $M$. Recall that for a square matrix $M$, we denote by $\mathop{\mathrm{Tr}}M$ and $\mathop{\mathrm{tr}}M$ the unnormalized and normalized trace, respectively.
We write $[n]:=\{1,\ldots,n\}$ for $n\in\mathbb{N}$. We denote by $\mathrm{P}_{2}([n])$ the set of all pairings of $[n]$ (that is, partitions each of whose elements has size two). Recall that given any $\pi\in\mathrm{P}_{2}([n])$, two pairs $\{i,k\},\{j,l\}\in\pi$ such that $i<j<k<l$ are said to form a crossing. A pairing that contains no crossing is said to be noncrossing, and we denote by $\mathrm{NC}_{2}([n])$ the set of all noncrossing pairings of $[n]$.
Finally, we will write $x\lesssim y$ to indicate that $x\leq Cy$ for a universal constant $C$.
2. The Hermitian case
The aim of this section is to investigate Hermitian random matrices with independent complex Gaussian entries. That is, we consider the following model.
Model 2.1.
$X$ is an $n\times n$ Hermitian matrix whose entries $X_{ij}=(X_{ji})^{*}$ are independent for $i\geq j$ with $X_{ij}\sim N_{\mathbb{C}}(0,b_{ij}^{2})$ for $i>j$ and $X_{ii}\sim N_{\mathbb{R}}(0,b_{ii}^{2})$, where $b_{ij}=b_{ji}\geq 0$ are arbitrary nonnegative scalars.
This setting admits particularly simple proofs, which will guide the more involved arguments required for real random matrices in the following sections.
Define the parameters
$\sigma^{2}:=\max_{i\leq n}\sum_{j\leq n}b_{ij}^{2},\qquad\qquad\sigma_{*}^{2}:=\max_{i,j\leq n}b_{ij}^{2}.$ (2.1)
Then we have the following extremum principle and tail bounds.
Theorem 2.2 (Extremum principle).
Define $X$ as in Model 2.1, and assume that $\sigma_{*}^{2}=1$ and that $\sigma^{2}\leq d\in\mathbb{N}$. Then we have
$\mathbf{E}[\mathop{\mathrm{tr}}X^{2p}]\leq\mathbf{E}[\mathop{\mathrm{tr}}Y^{2p}]$
for all $p\in\mathbb{N}$, where $Y$ is the $d\times d$ Hermitian matrix whose entries $Y_{ij}=(Y_{ji})^{*}$ are independent for $i\geq j$ with $Y_{ij}\sim N_{\mathbb{C}}(0,1)$ for $i>j$ and $Y_{ii}\sim N_{\mathbb{R}}(0,1)$.
Theorem 2.3 (Small deviations).
For $X$ as in Model 2.1, we have
$\mathbf{P}\big{[}\|X\|>2\sigma+4\sigma_{*}^{4/3}\sigma^{-1/3}t\big{]}\leq\frac{en\sigma_{*}^{2}}{\sigma^{2}}e^{-t^{3/2}}$
for every $0\leq t\leq\frac{\sigma^{4/3}}{\sigma_{*}^{4/3}}$.
Theorem 2.4 (Large deviations).
For $X$ as in Model 2.1, we have
$\mathbf{P}\big{[}\|X\|>2\sigma+\sigma_{*}(1+t)\big{]}\leq 2n\,e^{-t^{2}/2}$
for every $t\geq 0$.
The remainder of this section is devoted to the proofs of these results.
2.1. Extremum principle
In this subsection, we will use the following notation. Let $(g_{ij})_{i\geq j}$ be i.i.d. $N_{\mathbb{C}}(0,1)$ random variables, and define
$U_{ij}:=b_{ij}e_{i}e_{j}^{*},\qquad\qquad U_{ii}:=\tfrac{1}{\sqrt{2}}b_{ii}e_{i}e_{i}^{*}$
for $i>j$. Then we can represent the random matrix $X$ of Model 2.1 as
$X=\sum_{i\geq j}(g_{ij}U_{ij}+g_{ij}^{*}U_{ij}^{*}).$
We can compute the moments of $X$ as follows.
Lemma 2.5 (Wick formula).
For any $p\in\mathbb{N}$, we have
$\mathbf{E}[\mathop{\mathrm{tr}}X^{2p}]=\sum_{\pi\in\mathrm{P}_{2}([2p])}\sum_{(\bm{i},\bm{j},\bm{\varepsilon})\sim\pi}\mathop{\mathrm{tr}}U_{i_{1}j_{1}}^{\varepsilon_{1}}\cdots U_{i_{2p}j_{2p}}^{\varepsilon_{2p}}.$
Here $\bm{i}=(i_{1},\ldots,i_{2p})\in[n]^{2p}$, $\bm{j}=(j_{1},\ldots,j_{2p})\in[n]^{2p}$, $\bm{\varepsilon}=(\varepsilon_{1},\ldots,\varepsilon_{2p})\in\{1,*\}^{2p}$, and $(\bm{i},\bm{j},\bm{\varepsilon})\sim\pi$ denotes $i_{k}\geq j_{k}$ and $i_{k}=i_{l},j_{k}=j_{l},\varepsilon_{k}\neq\varepsilon_{l}$ for all $\{k,l\}\in\pi$.
Proof.
Clearly
$\mathbf{E}[\mathop{\mathrm{tr}}X^{2p}]=\sum_{i_{1}\geq j_{1},\ldots,i_{2p}\geq j_{2p}}\sum_{\varepsilon_{1},\ldots,\varepsilon_{2p}\in\{1,*\}}\mathbf{E}[g_{i_{1}j_{1}}^{\varepsilon_{1}}\cdots g_{i_{2p}j_{2p}}^{\varepsilon_{2p}}]\mathop{\mathrm{tr}}U_{i_{1}j_{1}}^{\varepsilon_{1}}\cdots U_{i_{2p}j_{2p}}^{\varepsilon_{2p}}.$
The conclusion follows as
$\mathbf{E}[g_{i_{1}j_{1}}^{\varepsilon_{1}}\cdots g_{i_{2p}j_{2p}}^{\varepsilon_{2p}}]=\sum_{\pi\in\mathrm{P}_{2}([2p])}\prod_{\{k,l\}\in\pi}1_{i_{k}=i_{l},j_{k}=j_{l},\varepsilon_{k}\neq\varepsilon_{l}}$
by the classical Wick formula [18, Theorem 22.3 and Remark 22.5]. ∎
Free probability theory suggests [18, p. 367] that the expression in Lemma 2.5 should be dominated by noncrossing pairings. The idea behind the proof of Theorem 2.2 is to consider separately the effect of noncrossing and crossing pairs on the terms in the sum in Lemma 2.5. We first consider noncrossing pairings.
Lemma 2.6.
For any $p\in\mathbb{N}$ and noncrossing pairing $\pi\in\mathrm{NC}_{2}([2p])$, we have
$\sum_{(\bm{i},\bm{j},\bm{\varepsilon})\sim\pi}\mathop{\mathrm{tr}}U_{i_{1}j_{1}}^{\varepsilon_{1}}\cdots U_{i_{2p}j_{2p}}^{\varepsilon_{2p}}\leq\sigma^{2p}.$
Moreover, equality holds when $b_{ij}=1$ for all $i,j$.
Proof.
Recall the elementary fact that any noncrossing pairing $\pi$ must contain a consecutive pair $\{k,k+1\}\in\pi$. (If not, choose $\{k,k+l\}\in\pi$ with minimal $l\geq 2$. Then $\{k+1,r\}\in\pi$ must satisfy $|k+1-r|\geq l$. In particular, $\{k,k+l\}$ and $\{k+1,r\}$ form a crossing, contradicting the assumption.)
By cyclic permutation of the trace, we may assume without loss of generality that $\{2p-1,2p\}\in\pi$. Then we can compute
$\sum_{(\bm{i},\bm{j},\bm{\varepsilon})\sim\pi}\mathop{\mathrm{tr}}U_{i_{1}j_{1}}^{\varepsilon_{1}}\cdots U_{i_{2p}j_{2p}}^{\varepsilon_{2p}}=\\ \sum_{(\bm{i},\bm{j},\bm{\varepsilon})\sim\pi\backslash\{\{2p-1,2p\}\}}\mathop{\mathrm{tr}}U_{i_{1}j_{1}}^{\varepsilon_{1}}\cdots U_{i_{2p-2}j_{2p-2}}^{\varepsilon_{2p-2}}\sum_{i\geq j}(U_{ij}U_{ij}^{*}+U_{ij}^{*}U_{ij}).$
Now note that
$\sum_{i\geq j}(U_{ij}U_{ij}^{*}+U_{ij}^{*}U_{ij})=\sum_{i\leq n}\bigg{(}\sum_{j\leq n}b_{ij}^{2}\bigg{)}e_{i}e_{i}^{*}.$
As all $U_{ij}$ have nonnegative entries, we have $\mathop{\mathrm{tr}}U_{i_{1}j_{1}}^{\varepsilon_{1}}\cdots U_{i_{2p-2}j_{2p-2}}^{\varepsilon_{2p-2}}e_{i}e_{i}^{*}\geq 0$. Thus
$\sum_{(\bm{i},\bm{j},\bm{\varepsilon})\sim\pi}\mathop{\mathrm{tr}}U_{i_{1}j_{1}}^{\varepsilon_{1}}\cdots U_{i_{2p}j_{2p}}^{\varepsilon_{2p}}\leq\sigma^{2}\sum_{(\bm{i},\bm{j},\bm{\varepsilon})\sim\pi\backslash\{\{2p-1,2p\}\}}\mathop{\mathrm{tr}}U_{i_{1}j_{1}}^{\varepsilon_{1}}\cdots U_{i_{2p-2}j_{2p-2}}^{\varepsilon_{2p-2}},$
with equality if $b_{ij}=1$ for all $i,j$. As $\pi\backslash\{\{2p-1,2p\}\}$ is again a noncrossing pairing, we can iterate this argument to conclude the proof. ∎
Next, we analyze the effect of a single crossing.
Lemma 2.7 (Crossing inequality).
Let $M_{1},\ldots,M_{4}$ be any $n\times n$ matrices with nonnegative entries. Then we have
$\sum_{i\geq j,k\geq l}\sum_{\varepsilon,\delta\in\{1,*\}}\mathop{\mathrm{tr}}U_{ij}^{\varepsilon}M_{1}U_{kl}^{\delta}M_{2}U_{ij}^{!\varepsilon}M_{3}U_{kl}^{!\delta}M_{4}\leq\sigma_{*}^{4}\mathop{\mathrm{tr}}M_{3}M_{2}M_{1}M_{4},$
with equality when $b_{ij}=1$ for all $i,j$. Here we denote !1:=*!1:=*! 1 := * and !*:=1!*:=1! * := 1.
Proof.
Substituting the definition of $U_{ij}$ into the left-hand side yields
$\sum_{i\geq j,k\geq l}\sum_{\varepsilon,\delta\in\{1,*\}}\mathop{\mathrm{tr}}U_{ij}^{\varepsilon}M_{1}U_{kl}^{\delta}M_{2}U_{ij}^{!\varepsilon}M_{3}U_{kl}^{!\delta}M_{4}=$
$\qquad\qquad\frac{1}{n}\sum_{i,j,k,l\in[n]}b_{ij}^{2}b_{kl}^{2}(M_{3})_{il}(M_{2})_{lj}(M_{1})_{jk}(M_{4})_{ki}$
by an explicit computation. The conclusion follows readily. ∎
We can now prove Theorem 2.2.
Proof of Theorem 2.2.
Fix a pairing $\pi\in\mathrm{P}_{2}([2p])$. Suppose that $\{1,l\},\{k,m\}$ form a crossing $1<k<l<m$. Then Lemma 2.7 yields
$\sum_{(\bm{i},\bm{j},\bm{\varepsilon})\sim\pi}\mathop{\mathrm{tr}}U_{i_{1}j_{1}}^{\varepsilon_{1}}\cdots U_{i_{2p}j_{2p}}^{\varepsilon_{2p}}\leq\sigma_{*}^{4}\sum_{(\bm{i},\bm{j},\bm{\varepsilon})\sim\pi^{\prime}}\mathop{\mathrm{tr}}U_{i_{1}j_{1}}^{\varepsilon_{1}}\cdots U_{i_{2p-4}j_{2p-4}}^{\varepsilon_{2p-4}}.$
where $\pi^{\prime}\in\mathrm{P}_{2}([2p-4])$ is obtained from $\pi$ by removing $\{1,l\},\{k,m\}$ and reordering the remaining elements as $l+1,\ldots,m-1,k+1,\ldots,l-1,2,\ldots,k-1,m+1,\ldots,2p$. (Here we used that any product of matrices $U_{ij}^{\varepsilon}$ has nonnegative elements.)
We can now iterate this procedure until we arrive at a final pairing $\pi^{\prime}$ that is noncrossing. More precisely, given any pairing $\pi$ that contains at least one crossing, we choose the smallest crossing in the lexicographic order and use cyclic permutation of the trace to make its smallest element equal to one. We then apply the above inequality. If $\pi^{\prime}$ still contains a crossing, we repeat this procedure until no crossings are left. Denote by $\ell(\pi)$ the number of times this process is repeated until we reach a noncrossing pairing. Then Lemmas 2.5 and 2.6 yield
$\mathbf{E}[\mathop{\mathrm{tr}}X^{2p}]\leq\sum_{\pi\in\mathrm{P}_{2}([2p])}\sigma^{2p-4\ell(\pi)}\sigma_{*}^{4\ell(\pi)}.$
Note that, by construction, the quantity $\ell(\pi)$ is uniquely determined by $\pi$ and does not depend on the random matrix $X$.
We now apply precisely the same argument to the random matrix $Y$. However, as every entry of $Y$ has equal variance, Lemmas 2.6 and 2.7 ensure that all the above inequalities become equalities in this case. In particular, we obtain
$\mathbf{E}[\mathop{\mathrm{tr}}Y^{2p}]=\sum_{\pi\in\mathrm{P}_{2}([2p])}d^{p-2\ell(\pi)}.$
As we assumed $\sigma^{2}\leq d$ and $\sigma_{*}^{2}=1$, the conclusion is immediate. ∎
Remark 2.8.
An identity of the above form for $\mathbf{E}[\mathop{\mathrm{tr}}Y^{2p}]$ is classical in random matrix theory, where it is known as the genus expansion [18, Theorem 22.12]. In particular, the precise combinatorial meaning of $\ell(\pi)$ can be understood in terms of the genus of the orientable surface obtained by gluing together the edges of a regular $2p$-gon corresponding to each pair of $\pi$. However, this combinatorial intepretation is completely irrelevant for our purposes: all we used is that the inequalities we apply to the nonhomogeneous matrix $X$ become equality for $Y$. For the real random matrices investigated in the following sections, the combinatorial structure is much more delicate while a comparison argument remains tractable.
2.2. Small deviations
We can now combine Theorem 2.2 with a method of Ledoux [15, §5.2] to obtain small deviations inequalities at the Tracy-Widom scale. The basis for this method is an accurate estimate on the $p$th moment of $Y$ for moderately large $p$. The following can be read off from [15, pp. 210–211].
Lemma 2.9 (Ledoux).
Define $Y$ as in Theorem 2.2. For all $p\in\mathbb{N}$, we have
$\mathbf{E}[\mathop{\mathrm{tr}}Y^{2p}]\leq\frac{1}{p^{3/2}\sqrt{\pi}}\bigg{(}4d+\frac{p(p-1)}{d}\bigg{)}^{p}.$
We obtain the following.
Proposition 2.10.
For $X$ as in Model 2.1 and $0\leq\varepsilon\leq 1$, we have
$\mathbf{P}\big{[}\|X\|>2\sqrt{\sigma^{2}+\sigma_{*}^{2}}\,(1+\varepsilon)\big{]}\leq\frac{en\sigma_{*}^{2}}{\sigma^{2}}e^{-\frac{\sigma^{2}}{\sigma_{*}^{2}}\varepsilon^{3/2}}.$
Proof.
Suppose first that $\sigma_{*}=1$, and let $d=\lceil\sigma^{2}\rceil$. Using Markov’s inequality, $\|X\|^{2p}\leq n\mathop{\mathrm{tr}}X^{2p}$, and Theorem 2.2, we obtain
$\mathbf{P}\big{[}\|X\|>2\sqrt{d}\,(1+\varepsilon)\big{]}\leq\frac{n}{(4d)^{p}}\frac{\mathbf{E}[\mathop{\mathrm{tr}}Y^{2p}]}{(1+\varepsilon)^{2p}}$
for all $p\in\mathbb{N}$. Applying Lemma 2.9 yields
$\mathbf{P}\big{[}\|X\|>2\sqrt{d}\,(1+\varepsilon)\big{]}\leq\frac{1}{\sqrt{\pi}}\frac{n}{p^{3/2}}e^{-\varepsilon p\log 4+p^{2}(p-1)/4d^{2}}$
using that $(1+\varepsilon)^{-2p}\leq 4^{-\varepsilon p}$ for $0<\varepsilon\leq 1$. Choosing $p=\lceil d\sqrt{2\varepsilon}\rceil$ yields
$\mathbf{P}\big{[}\|X\|>2\sqrt{d}\,(1+\varepsilon)\big{]}\leq 1.3\,\frac{n}{d}\frac{1}{(d\varepsilon^{3/2})^{1/2}}e^{-d\varepsilon^{3/2}}$
using $p^{2}(p-1)\leq(d\sqrt{2\varepsilon}+1)^{2}d\sqrt{2\varepsilon}\leq 2\sqrt{2}\,d^{3}\varepsilon^{3/2}+4d^{2}+d\sqrt{2}$ and $\sqrt{2}\log 4-\frac{1}{\sqrt{2}}\geq 1$.
We now consider two cases. First, if $\sigma^{2}\varepsilon^{3/2}\geq 1$, we can estimate
$\mathbf{P}\big{[}\|X\|>2\sqrt{d}\,(1+\varepsilon)\big{]}\leq\frac{en}{\sigma^{2}}e^{-\sigma^{2}\varepsilon^{3/2}}$
using $e\geq 1.3$ and $d\geq\sigma^{2}$. On the other hand, for $\sigma^{2}\varepsilon^{3/2}<1$, we have
$\mathbf{P}\big{[}\|X\|>2\sqrt{d}\,(1+\varepsilon)\big{]}\leq 1\leq\frac{en}{\sigma^{2}}e^{-\sigma^{2}\varepsilon^{3/2}}$
as $\sigma^{2}\leq n\sigma_{*}^{2}=n$. Combining these bounds, we obtain
$\mathbf{P}\big{[}\|X\|>2\sqrt{\sigma^{2}+1}\,(1+\varepsilon)\big{]}\leq\frac{en}{\sigma^{2}}e^{-\sigma^{2}\varepsilon^{3/2}}.$
where we used $d\leq\sigma^{2}+1$. This concludes the proof when $\sigma_{*}=1$. For general $\sigma_{*}$, it suffices to apply the above bound to the random matrix $\frac{X}{\sigma_{*}}$. ∎
Theorem 2.3 follows readily.
Proof of Theorem 2.3.
Proposition 2.10 implies
$\mathbf{P}\big{[}\|X\|>2\sqrt{\sigma^{2}+\sigma_{*}^{2}}+2\sqrt{\sigma^{2}+\sigma_{*}^{2}}\,\sigma_{*}^{4/3}\sigma^{-4/3}t\big{]}\leq\frac{en\sigma_{*}^{2}}{\sigma^{2}}e^{-t^{3/2}}$
by setting $\varepsilon=t\sigma_{*}^{4/3}\sigma^{-4/3}$. We now bound
$\sqrt{\sigma^{2}+\sigma_{*}^{2}}\leq\sigma+\frac{\sigma_{*}^{2}}{2\sigma}\leq\frac{3}{2}\sigma$
to obtain
$\mathbf{P}\big{[}\|X\|\geq 2\sigma+\tfrac{\sigma_{*}^{2}}{\sigma}+3\sigma_{*}^{4/3}\sigma^{-1/3}t\big{]}\leq\frac{en\sigma_{*}^{2}}{\sigma^{2}}e^{-t^{3/2}}.$
This readily implies the conclusion for $t\geq\sigma_{*}^{2/3}\sigma^{-2/3}$. On the other hand, as
$1\leq\frac{n\sigma_{*}^{2}}{\sigma^{2}}\leq\frac{en\sigma_{*}^{2}}{\sigma^{2}}e^{-t^{3/2}}$
for $t<\sigma_{*}^{2/3}\sigma^{-2/3}\leq 1$, the conclusion holds trivially in this case. ∎
2.3. Large deviations
The moment estimate of Lemma 2.9 is not accurate for large $p$: indeed, this estimate yields a subexponential bound $\mathbf{E}[\mathop{\mathrm{tr}}Y^{2p}]^{1/2p}=O(p)$ as $p\to\infty$, while the large deviations of the norms of Gaussian random matrices are in fact subgaussian. To obtain a large deviation inequality, we will instead employ the following estimate of Haagerup and Thorbjørnsen [10, Eq. (3.5)].
Lemma 2.11 (Haagerup-Thorbjørnsen).
Let $Y$ be as in Theorem 2.2. For all $t\geq 0$
$\mathbf{E}[\mathop{\mathrm{tr}}e^{tY}]\leq e^{2\sqrt{d}\,t+t^{2}/2}.$
This yields the following.
Proposition 2.12.
For $X$ as in Model 2.1 and $\varepsilon\geq 0$, we have
$\mathbf{P}\big{[}\|X\|>2\sqrt{\sigma^{2}+\sigma_{*}^{2}}+\sigma_{*}\varepsilon\big{]}\leq 2n\,e^{-\varepsilon^{2}/2}.$
Proof.
Suppose first that $\sigma_{*}=1$, and let $d=\lceil\sigma^{2}\rceil$. Note that all odd moments of our random matrices vanish $\mathbf{E}[\mathop{\mathrm{tr}}X^{2p+1}]=\mathbf{E}[\mathop{\mathrm{tr}}Y^{2p+1}]=0$ by the symmetry of the Gaussian distribution. We can therefore estimate
$\mathbf{E}[\mathop{\mathrm{tr}}e^{tX}]\leq\mathbf{E}[\mathop{\mathrm{tr}}e^{tY}]$
for $t\geq 0$ by Taylor expanding the exponential function and applying Theorem 2.2 to the terms of even degree. As $e^{t\|X\|}\leq n\mathop{\mathrm{tr}}e^{tX}+n\mathop{\mathrm{tr}}e^{-tX}$, we obtain
$\mathbf{E}[e^{t\|X\|}]\leq 2n\,e^{2\sqrt{d}\,t+t^{2}/2}$
using Lemma 2.11. By Markov’s inequality
$\mathbf{P}\big{[}\|X\|>2\sqrt{d}+\varepsilon\big{]}\leq\frac{\mathbf{E}[e^{t\|X\|}]}{e^{(2\sqrt{d}+\varepsilon)t}}\leq 2n\,e^{-\varepsilon t+t^{2}/2}.$
Optimizing over $t\geq 0$ yields
$\mathbf{P}\big{[}\|X\|>2\sqrt{\sigma^{2}+1}+\varepsilon\big{]}\leq 2n\,e^{-\varepsilon^{2}/2}$
for all $\varepsilon\geq 0$, where we used $d\leq\sigma^{2}+1$. This concludes the proof for $\sigma_{*}=1$. For general $\sigma_{*}$, it suffices to apply the above bound to the random matrix $\frac{X}{\sigma_{*}}$. ∎
Theorem 2.4 follows readily.
Proof of Theorem 2.4.
The conclusion follows immediately from Proposition 2.12 using that $\sqrt{\sigma^{2}+\sigma_{*}^{2}}\leq\sigma+\frac{\sigma_{*}^{2}}{2\sigma}\leq\sigma+\frac{\sigma_{*}}{2}$. ∎
3. The symmetric case
We now turn to the case of (real) symmetric matrices. The following model is the real symmetric analogue of Model 1.3.
Model 3.1.
$X$ is an $n\times n$ real symmetric matrix with $X_{ij}=b_{ij}\xi_{ij}$ for $i\neq j$ and $X_{ii}=\sqrt{2}\,b_{ii}\xi_{ii}$. Here $\xi_{ij}=\xi_{ji}$ are independent symmetrically distributed real random variables for $i\geq j$ with $\mathbf{E}[\xi_{ij}^{2p}]\leq\mathbf{E}[g^{2p}]$ for all $p\in\mathbb{N}$ (here $g\sim N(0,1)$), and $b_{ij}=b_{ji}\geq 0$ are arbitrary nonnegative scalars.
Remark 3.2.
The slightly different scaling of the off-diagonal and diagonal entries ensures that $X$ is GOE (a real symmetric random matrix whose law is invariant under orthogonal conjugation) when $b_{ij}=1$ for all $i,j$ and $\xi_{ij}$ are Gaussian.
In this setting, the parameters $\sigma$ and $\sigma_{*}$ are defined as in (2.1). The main results of this section are the following extremum principle and tail bounds.
Theorem 3.3 (Extremum principle).
Define $X$ as in Model 3.1, and assume that $\sigma_{*}^{2}=1$ and that $\sigma^{2}\leq d\in\mathbb{N}$. Then we have
$\mathbf{E}[\mathop{\mathrm{tr}}X^{2p}]\leq\mathbf{E}[\mathop{\mathrm{tr}}Y^{2p}]$
for all $p\in\mathbb{N}$, where $Y$ is the $d\times d$ real symmetric matrix whose entries $Y_{ij}=Y_{ji}$ are independent for $i\geq j$ with $Y_{ij}\sim N(0,1)$ for $i>j$ and $Y_{ii}\sim N(0,2)$.
Theorem 3.4 (Small deviations).
For $X$ as in Model 3.1, we have
$\mathbf{P}\big{[}\|X\|>2\sigma+\sigma_{*}^{4/3}\sigma^{-1/3}t\big{]}\leq\frac{n\sigma_{*}^{2}}{C\sigma^{2}}e^{-Ct^{3/2}}$
for every $0\leq t\leq\frac{\sigma^{4/3}}{\sigma_{*}^{4/3}}$, where $C$ is a universal constant.
Theorem 3.5 (Large deviations).
For $X$ as in Model 3.1, we have
$\mathbf{P}\big{[}\|X\|>2\sigma+\sigma_{*}(1+t)\big{]}\leq 2n\,e^{-t^{2}/4}$
for every $t\geq 0$.
The remainder of this section is devoted to the proofs of these results.
Remark 3.6.
The scale of the fluctuations in the tail bound of Theorem 3.5 is $\sqrt{2}\sigma_{*}$, while the scale of the fluctuations in Theorem 2.4 is only $\sigma_{*}$. The larger tail in the real symmetric case is necessary, however: for example, Theorem 3.5 implies that
[…]
as $n\to\infty$, while we have
[…]
whenever $b_{ii}=\sigma_{*}$ and $\xi_{ii}\sim N(0,1)$ for all $i$ (as then $X_{ii}=\sqrt{2}\sigma_{*}\xi_{ii}$). Moreover, it follows from [17, Corollary 3.2] that the tail bounds of both Theorems 2.4 and 3.5 cannot be improved in general for $t\to\infty$ (even in models where $b_{ii}=0$ for all $i$).
3.1. Extremum principle
Let $(g_{ij})_{i\geq j}$ be i.i.d. $N(0,1)$ random variables. We first show that there is no loss in assuming that $\xi_{ij}=g_{ij}$ in Model 3.1.
Lemma 3.7.
Let $X$ be as in Model 3.1, and let $\tilde{X}$ be defined as $X$ where $\xi_{ij}$ is replaced by $g_{ij}$. Then $\mathbf{E}[\mathop{\mathrm{tr}}X^{2p}]\leq\mathbf{E}[\mathop{\mathrm{tr}}\tilde{X}^{2p}]$ for all $p\in\mathbb{N}$.
Proof.
Note that $\mathop{\mathrm{tr}}X^{2p}$ is a polynomial of $(X_{ij})_{i\geq j}$ with nonnegative coefficients, and that the expectation of each monomial is either zero or a product of terms of the form $\mathbf{E}[X_{ij}^{2k}]$ ($k\in\mathbb{N}$) as $X_{ij}$ are symmetrically distributed and independent. The conclusion follows from the assumption that $\mathbf{E}[\xi_{ij}^{2k}]\leq\mathbf{E}[g_{ij}^{2k}]$ for all $k\in\mathbb{N}$. ∎
We will therefore assume in the remainder of the proof of Theorem 3.3 that $X$ is defined as in Model 3.1 with $\xi_{ij}=g_{ij}$. Then we define
$H_{ij}:=b_{ij}(e_{i}e_{j}^{*}+e_{j}e_{i}^{*}),\qquad\qquad H_{ii}:=\sqrt{2}\,b_{ii}e_{i}e_{i}^{*}$
for $i>j$, and represent the random matrix $X$ as
$X=\sum_{i\geq j}g_{ij}H_{ij}.$
We can now compute the moments of $X$ as follows.
Lemma 3.8 (Wick formula).
For any $p\in\mathbb{N}$, we have
$\mathbf{E}[\mathop{\mathrm{tr}}X^{2p}]=\sum_{\pi\in\mathrm{P}_{2}([2p])}\mathop{\mathrm{tr}}H(\pi)$
with
$H(\pi):=\sum_{(\bm{i},\bm{j})\sim\pi}H_{i_{1}j_{1}}\cdots H_{i_{2p}j_{2p}}.$
Here $\bm{i},\bm{j}\in[n]^{2p}$, and $(\bm{i},\bm{j})\sim\pi$ denotes $i_{k}\geq j_{k}$ and $i_{k}=i_{l},j_{k}=j_{l}$ for $\{k,l\}\in\pi$.
Proof.
The proof is identical to that of Lemma 2.5. ∎
The challenge in the real case is that the effect of a crossing is much more complicated than in the complex case (Lemma 2.7): we will have to distinguish in the analysis between two different types of crossings that contribute in a different way to the Wick formula. In order to do so efficiently, it will turn out to be necessary to work with the following upper bound on the trace moments.
Lemma 3.9.
For any $k\in\mathbb{N}$ and $\pi\in\mathrm{P}_{2}([2k])$, define
$C(\pi):=\max_{r}(H(\pi))_{rr}.$
Then
$\mathbf{E}[\mathop{\mathrm{tr}}X^{2p}]\leq\sum_{\pi\in\mathrm{P}_{2}([2p])}C(\pi).$
Moreover, when $b_{ij}=1$ for all $i,j$, the matrix $H(\pi)$ is a multiple of the identity matrix for every $\pi$ and thus the above inequality holds with equality.
Proof.
The inequality is immediate from Lemma 3.8. It remains to show that $H(\pi)$ is a multiple of the identity matrix when $b_{ij}=1$ for all $i,j$. To this end, note that we may write $H(\pi)=\mathbf{E}[X_{1}\cdots X_{2k}]$, where $X_{k}=X_{l}$ is an independent copy of $X$ for each $\{k,l\}\in\pi$. When $b_{ij}\equiv 1$, $X$ is a GOE matrix and its distribution is therefore orthogonally invariant. Thus $O\mathbf{E}[X_{1}\cdots X_{2k}]O^{*}=\mathbf{E}[OX_{1}O^{*}\cdots OX_{2k}O^{*}]=\mathbf{E}[X_{1}\cdots X_{2k}]$ for every $O\in O(n)$, from which the conclusion follows directly. ∎
We will now analyze the quantity $C(\pi)$. We first consider noncrossing pairings.
Lemma 3.10.
For any $p\in\mathbb{N}$ and noncrossing pairing $\pi\in\mathrm{NC}_{2}([2p])$, we have
$C(\pi)\leq\tilde{\sigma}^{2p},\qquad\quad\tilde{\sigma}^{2}:=\max_{i}\bigg{(}\sum_{j}b_{ij}^{2}+b_{ii}^{2}\bigg{)}.$
Moreover, equality holds when $b_{ij}=1$ for all $i,j$.
Proof.
As in the proof of Lemma 2.6, there exists $\{k,k+1\}\in\pi$. Then
$C(\pi)=\max_{r}\sum_{(\bm{i},\bm{j})\sim\pi\backslash\{\{k,k+1\}\}}(H_{i_{1}j_{1}}\cdots H_{i_{k-1}j_{k-1}}\Sigma H_{i_{k+2}j_{k+2}}\cdots H_{i_{2p}j_{2p}})_{rr},$
where
$\Sigma:=\sum_{i\geq j}H_{ij}^{2}=\sum_{i}\bigg{(}\sum_{j}b_{ij}^{2}+b_{ii}^{2}\bigg{)}e_{i}e_{i}^{*}.$
But note that $(H_{i_{1}j_{1}}\cdots H_{i_{k-1}j_{k-1}}e_{i}e_{i}^{*}H_{i_{k+2}j_{k+2}}\cdots H_{i_{2k}j_{2k}})_{rr}\geq 0$ as all $H_{ij}$ have nonnegative entries. We can therefore readily estimate
$C(\pi)\leq\tilde{\sigma}^{2}\,C(\pi\backslash\{\{k,k+1\}\}),$
with equality if $b_{ij}=1$ for all $i,j$. As $\pi\backslash\{\{k,k+1\}\}$ is again noncrossing, we can iterate this argument to conclude the proof. ∎
We now consider the effect of a crossing. We will need the following identities, which follow from somewhat tedious but straightforward computations.
Lemma 3.11 (Crossing identities).
The following hold.
  1. a.
    For any matrices $M_{1},M_{2},M_{3}$, we have
    $\sum_{i\geq j}\sum_{k\geq l}H_{ij}M_{1}H_{kl}M_{2}H_{ij}M_{3}H_{kl}=\sum_{i,j,k,l}b_{ij}^{2}b_{kl}^{2}\big{\{}(M_{1})_{jk}(M_{2})_{lj}(M_{3})_{il}+\mbox{}\\ (M_{1})_{jl}(M_{2})_{kj}(M_{3})_{il}+(M_{1})_{jk}(M_{2})_{li}(M_{3})_{jl}+(M_{1})_{jl}(M_{2})_{ki}(M_{3})_{jl}\big{\}}e_{i}e_{k}^{*}.$
  2. b.
    For any matrix $M$, we have
    $\sum_{i\geq j}\mathrm{Tr}[MH_{ij}]H_{ij}=\sum_{i,j}b_{ij}^{2}(M_{ij}+M_{ji})e_{i}e_{j}^{*}.$
Proof.
We readily compute that for any matrix $M$, we have
$\sum_{i\geq j}H_{ij}MH_{ij}=\sum_{i,j}b_{ij}^{2}(M_{jj}e_{i}e_{i}^{*}+M_{ji}e_{i}e_{j}^{*}).$
The first part of the statement follows by applying this identity twice. The second part of the statement is again a straightforward computation. ∎
To proceed, we must distinguish between two types of crossings that play a fundamentally different role in the argument.
Definition 3.12.
Let $\pi\in\mathrm{P}_{2}([2p])$, and fix a crossing $\{i,k\},\{j,l\}\in\pi$ such that $i<j<k<l$. Then this crossing is said to be of type I if there exists $\{a,b\}\in\pi$ with $a\in(i,j)\cup(k,l)$ and $b\not\in(i,j)\cup(k,l)$, an is said to be of type II otherwise.
We consider each type of crossing in turn.
Lemma 3.13 (Type I crossing).
Let $p\in\mathbb{N}$, $\pi\in\mathrm{P}_{2}([2p])$, $\{a,c\},\{b,d\},\{e,f\}\in\pi$ with $a<b<c<d$ and $e\in(a,b)\cup(c,d)$, $f\not\in(a,b)\cup(c,d)$. Then there exist pairings $\pi_{1},\pi_{2},\pi_{3}\in\mathrm{P}_{2}([2p-4])$ and $\pi_{4},\pi_{5}\in\mathrm{P}_{2}([2p-6])$, whose definition depends only on $\pi,a,b,c,d,e,f$ (and not on the matrix $X$), so that [^2] Here $C(\pi_{4})=C(\varnothing):=1$ if $02p-6=$ 0. The analogous convention will be used in the sequel.
$C(\pi)\leq\sigma_{*}^{4}\{C(\pi_{1})+C(\pi_{2})+C(\pi_{3})\}+\sigma_{*}^{6}\,\{C(\pi_{4})+C(\pi_{5})\}.$
Moreover, equality holds when $b_{ij}=1$ for all $i,j$.
Proof.
For any $M_{1},M_{2},M_{3}$ with nonnegative entries, Lemma 3.11(a) yields
$\sum_{i\geq j}\sum_{k\geq l}(H_{ij}M_{1}H_{kl}M_{2}H_{ij}M_{3}H_{kl})_{rs}\\ \leq\sigma_{*}^{4}\big{(}M_{3}M_{2}M_{1}+M_{3}M_{1}^{*}M_{2}^{*}+M_{2}^{*}M_{2}^{*}M_{1}+\mathrm{Tr}[M_{1}^{*}M_{3}]M_{2}^{*}\big{)}_{rs}$
with equality if $b_{ij}=1$ for all $i,j$. We can therefore write
$C(\pi)\leq\sigma_{*}^{4}\max_{r}\sum_{(\bm{i},\bm{j})\sim\pi\backslash\{\{a,c\},\{b,d\}\}}\big{\{}(M_{0}M_{3}M_{2}M_{1}M_{4})_{rr}+(M_{0}M_{3}M_{1}^{*}M_{2}^{*}M_{4})_{rr}\\ +(M_{0}M_{2}^{*}M_{2}^{*}M_{1}M_{4})_{rr}+\mathrm{Tr}[M_{1}^{*}M_{3}](M_{0}M_{2}^{*}M_{4})_{rr}\big{\}}$
with equality if $b_{ij}=1$ for all $i,j$, where
$M_{0}$ $:=H_{i_{1}j_{1}}\cdots H_{i_{a-1}j_{a-1}},$
$M_{1}$ $:=H_{i_{a+1}j_{a+1}}\cdots H_{i_{b-1}j_{b-1}},$
$M_{2}$ $:=H_{i_{b+1}j_{b+1}}\cdots H_{i_{c-1}j_{c-1}},$
$M_{3}$ $:=H_{i_{c+1}j_{c+1}}\cdots H_{i_{d-1}j_{d-1}},$
$M_{4}$ $:=H_{i_{d+1}j_{d+1}}\cdots H_{i_{2p}j_{2p}}.$
For simplicity, we consider in the remainder of the proof the case that $c<e<d$ and $f>d$; the proof of the five other possible cases is completely analogous. Then we can use Lemma 3.11(b) to estimate
$\sum_{(\bm{i},\bm{j})\sim\pi\backslash\{\{a,c\},\{b,d\}\}}\mathrm{Tr}[M_{1}^{*}M_{3}](M_{0}M_{2}^{*}M_{4})_{rr}$
$=\sum_{(\bm{i},\bm{j})\sim\pi\backslash\{\{a,c\},\{b,d\},\{e,f\}\}}\sum_{i\geq j}\mathrm{Tr}[M_{1}^{*}M_{3}^{-}H_{ij}M_{3}^{+}](M_{0}M_{2}^{*}M_{4}^{-}H_{ij}M_{4}^{+})_{rr}$
$\leq\sigma_{*}^{2}\sum_{(\bm{i},\bm{j})\sim\pi\backslash\{\{a,c\},\{b,d\},\{e,f\}\}}\big{\{}(M_{0}M_{2}^{*}M_{4}^{-}M_{3}^{+}M_{1}^{*}M_{3}^{-}M_{4}^{+})_{rr}$
$\qquad\qquad\qquad\qquad\qquad\qquad\qquad+(M_{0}M_{2}^{*}M_{4}^{-}M_{3}^{-*}M_{1}M_{3}^{+*}M_{4}^{+})_{rr}\big{\}}$
with equality if $b_{ij}=1$ for all $i,j$, where
$M_{3}^{-}$ $:=H_{i_{c+1}j_{c+1}}\cdots H_{i_{e-1}j_{e-1}},$
$M_{3}^{+}$ $:=H_{i_{e+1}j_{e+1}}\cdots H_{i_{d-1}j_{d-1}},$
$M_{4}^{-}$ $:=H_{i_{d+1}j_{d+1}}\cdots H_{i_{f-1}j_{f-1}},$
$M_{4}^{+}$ $:=H_{i_{f+1}j_{f+1}}\cdots H_{i_{2p}j_{2p}}.$
From the above, we can readily read off the existence of $\pi_{1},\pi_{2},\pi_{3}\in\mathrm{P}_{2}([2p-4])$ and $\pi_{4},\pi_{5}\in\mathrm{P}_{2}([2p-6])$, whose definition depends only on $\pi,a,b,c,d,e,f$, so that
$C(\pi)\leq\max_{r}\big{\{}\sigma_{*}^{4}\big{(}H(\pi_{1})+H(\pi_{2})+H(\pi_{3})\big{)}_{rr}+\sigma_{*}^{6}\big{(}H(\pi_{4})+H(\pi_{5})\big{)}_{rr}\big{\}}$
with equality if $b_{ij}=1$ for all $i,j$, where $H(\pi)$ was defined in Lemma 3.8. The inequality in the statement follows immediately by using $(H(\pi))_{rr}\leq C(\pi)$ for all $r,\pi$. On the other hand, when $b_{ij}=1$ for all $i,j$, Lemma 3.9 yields $H(\pi)=C(\pi)\mathbf{1}$ for all $\pi$, so that the inequality in the statement holds with equality. ∎
When Lemma 3.11(a) is applied as in the proof of Lemma 3.13, it creates a term that involves an unnormalized trace. For crossings of type I, we can subsequently apply Lemma 3.11(b) to eliminate the trace and regain a term of the form $H(\pi)$. This is not possible, however, for crossings of type II, which would cause the trace to factor out of the expression. As this trace is unnormalized, that would ultimately yield a dimension-dependent factor in the final bound.
To avoid this, we must apply Lemma 3.11(a) in a different manner to crossings of type II. It is here that we rely on the fact that we work with the quantity $C(\pi)$, rather than the smaller quantity $\mathop{\mathrm{tr}}H(\pi)$ that appears in the Wick formula itself. (This is in contrast to the statement and proof of Lemma 3.13, which would work equally well if we replace $C(\pi)$ by $\mathop{\mathrm{tr}}H(\pi)$ throughout.)
Lemma 3.14 (Type II crossing).
Let $p\in\mathbb{N}$, $\pi\in\mathrm{P}_{2}([2p])$, and let $\{a,c\},\{b,d\}\in\pi$ with $a<b<c<d$ be a crossing of type II. Then there exist pairings $\pi_{1},\pi_{2},\pi_{3}\in\mathrm{P}_{2}([2p-4])$ and $\pi_{4}\in\mathrm{P}_{2}([b-a+d-c-2])$, $\pi_{5}\in\mathrm{P}_{2}([2p-2-b+a-d+c])$, whose definition depends only on $\pi,a,b,c,d$ (and not on the matrix $X$), so that
$C(\pi)\leq\sigma_{*}^{4}\{C(\pi_{1})+C(\pi_{2})+C(\pi_{3})\}+\sigma^{2}\sigma_{*}^{2}\,C(\pi_{4})\,C(\pi_{5}).$
Moreover, equality holds when $b_{ij}=1$ for all $i,j$.
Proof.
Define $M_{0},\ldots,M_{4}$ (which depend on $\bm{i},\bm{j}$) and $\pi_{1},\pi_{2},\pi_{3}$ as in the proof of Lemma 3.13. Then we can estimate using Lemma 3.11(a)
$C(\pi)\leq\max_{r}\bigg{\{}\sigma_{*}^{4}\big{(}H(\pi_{1})+H(\pi_{2})+H(\pi_{3})\big{)}_{rr}+\\ \sigma_{*}^{2}\sum_{(\bm{i},\bm{j})\sim\pi\backslash\{\{a,c\},\{b,d\}\}}\sum_{k,l}b_{kl}^{2}(M_{0}M_{2}^{*})_{rk}(M_{1}^{*}M_{3})_{ll}(M_{4})_{kr}\bigg{\}},$
with equality if $b_{ij}=1$ for all $i,j$. Note that we applied Lemma 3.11(a) here exactly as in the proof of Lemma 3.13, except that we only upper bounded $b_{ij}^{2}\leq\sigma_{*}^{2}$ in the last term of Lemma 3.11(a) and avoided upper bounding $b_{kl}^{2}$ as well.
To proceed more efficiently in the present setting, we will now use the special structure of type II crossings. By the type II assumption, we can decompose $\pi\backslash\{\{a,c\},\{b,d\}\}=\pi_{I}\cup\pi_{J}$ where $\pi_{I}\in\mathrm{P}_{2}(I)$, $\pi_{J}\in\mathrm{P}_{2}(J)$ with
$I:=(a,b)\cup(c,d),\qquad\quad J:=(1,a)\cup(b,c)\cup(d,2p).$
Denote by $\bm{i}_{I}:=(i_{s})_{s\in I}$, with the analogous notation for $\bm{j}$ and $J$. Then
$\sum_{(\bm{i},\bm{j})\sim\pi\backslash\{\{a,c\},\{b,d\}\}}\sum_{k,l}b_{kl}^{2}(M_{0}M_{2}^{*})_{rk}(M_{1}^{*}M_{3})_{ll}(M_{4})_{kr}$
$\qquad=\sum_{k}\Bigg{(}\sum_{(\bm{i}_{J},\bm{j}_{J})\sim\pi_{J}}(M_{0}M_{2}^{*})_{rk}(M_{4})_{kr}\Bigg{)}\Bigg{(}\sum_{l}b_{kl}^{2}\sum_{(\bm{i}_{I},\bm{j}_{I})\sim\pi_{I}}(M_{1}^{*}M_{3})_{ll}\Bigg{)}$
$\qquad\leq\sigma^{2}\Bigg{(}\sum_{(\bm{i}_{J},\bm{j}_{J})\sim\pi_{J}}(M_{0}M_{2}^{*}M_{4})_{rr}\Bigg{)}\max_{l}\Bigg{(}\sum_{(\bm{i}_{I},\bm{j}_{I})\sim\pi_{I}}(M_{1}^{*}M_{3})_{ll}\Bigg{)},$
with equality if $b_{ij}=1$ for all $i,j$ (for the equality case, we used the second part of Lemma 3.9). From the above, we can readily read off the existence of pairings $\pi_{4}\in\mathrm{P}_{2}([b-a+d-c-2])$ and $\pi_{5}\in\mathrm{P}_{2}([2p-2-b+a-d+c])$, whose definition depends only on $\pi,a,b,c,d$, such that
$C(\pi)\leq\max_{r}\big{\{}\sigma_{*}^{4}\big{(}H(\pi_{1})+H(\pi_{2})+H(\pi_{3})\big{)}_{rr}+\sigma^{2}\sigma_{*}^{2}C(\pi_{4})(H(\pi_{5}))_{rr}\big{\}}$
with equality if $b_{ij}=1$ for all $i,j$. The conclusion now follows by precisely the same argument as at the end of the proof of Lemma 3.13. ∎
We can now conclude the proof of Theorem 3.3.
Proof of Theorem 3.3.
We first estimate $\mathbf{E}[\mathop{\mathrm{tr}}X^{2p}]$ as in Lemma 3.9. We then repeatedly apply the following to each quantity $C(\pi)$ in the resulting expression:
  1. $\bullet$
    If $\pi$ contains a crossing, we apply either Lemma 3.13 or Lemma 3.14 to the smallest crossing in the lexicographic order, depending on whether that crossing is of type I or type II, respectively. In the former case, we choose the smallest pair $\{e,f\}$ in the lexicographic order that satisfies the assumption of Lemma 3.13.
  2. $\bullet$
    If $\pi$ is noncrossing, we apply Lemma 3.10.
This algorithm gives rise to an inequality of the form
$\mathbf{E}[\mathop{\mathrm{tr}}X^{2p}]\leq\sum_{k,l\geq 0:k+l\leq p}\varkappa_{p}(k,l)\tilde{\sigma}^{2k}\sigma^{2l}\sigma_{*}^{2p-2k-2l},$
where the coefficients $\varkappa_{p}(k,l)\in\mathbb{Z}_{+}$ are independent of the matrix $X$.
Moreover, as all the inequalities used above become equalities when $b_{ij}=1$ for all $i,j$, we can apply the same algorithm to compute
$\mathbf{E}[\mathop{\mathrm{tr}}Y^{2p}]=\sum_{k,l\geq 0:k+l\leq p}\varkappa_{p}(k,l)(d+1)^{k}d^{l}.$
The conclusion readily follows from the assumptions that $\sigma_{*}=1$ and $\sigma^{2}\leq d$, and as we can estimate $\tilde{\sigma}^{2}\leq\sigma^{2}+1\leq d+1$ using $\sigma_{*}=1$. ∎
3.2. Small deviations
With the extremum principle in hand, we can now repeat the arguments of section 2.2 in the symmetric case. The main difference is that now $Y$ is a GOE matrix rather than a GUE matrix. However, a suitable moment estimate for GOE matrices was also obtained by Ledoux [16, Theorem 8].
Lemma 3.15 (Ledoux).
Define $Y$ as in Theorem 3.3. Then for $p\geq d^{2/3}$, we have
$\mathbf{E}[\mathop{\mathrm{tr}}Y^{2p}]\lesssim\frac{1}{d}(4d)^{p}\bigg{(}1+\frac{p^{2}}{d^{2}}\bigg{)}^{2p}.$
This yields the following.
Proposition 3.16.
For $X$ as in Model 3.1, we have
$\mathbf{P}\big{[}\|X\|>2\sqrt{\sigma^{2}+\sigma_{*}^{2}}\,(1+\varepsilon)\big{]}\leq\frac{n\sigma_{*}^{2}}{C\sigma^{2}}e^{-\frac{1}{4}\frac{\sigma^{2}}{\sigma_{*}^{2}}\varepsilon^{3/2}}$
for all $0\leq\varepsilon\leq 1$, where $C$ is a universal constant.
Proof.
Suppose first that $\sigma_{*}=1$, and let $d=\lceil\sigma^{2}\rceil$. Using Markov’s inequality, $\|X\|^{2p}\leq n\mathop{\mathrm{tr}}X^{2p}$, Theorem 3.3, and Lemma 3.15, we obtain
$\mathbf{P}\big{[}\|X\|>2\sqrt{d}\,(1+\varepsilon)\big{]}\lesssim\frac{n}{d}(1+\varepsilon)^{-2p}\bigg{(}1+\frac{p^{2}}{d^{2}}\bigg{)}^{2p}$
for all $p\geq d^{2/3}$. In particular, $(1+\varepsilon)^{-2p}\leq e^{-\varepsilon p}$ for $0<\varepsilon\leq 1$ and $1+x\leq e^{x}$ yield
$\mathbf{P}\big{[}\|X\|>2\sqrt{d}\,(1+\varepsilon)\big{]}\lesssim\frac{n}{d}\,e^{-\varepsilon p+2p^{3}/d^{2}}$
for all $p\geq d^{2/3}$.
We now consider two cases. If $\varepsilon\geq 16d^{-2/3}$, choose $p=\lfloor\frac{1}{2}d\sqrt{\varepsilon}\rfloor\geq d^{2/3}$ to obtain
$\mathbf{P}\big{[}\|X\|>2\sqrt{d}\,(1+\varepsilon)\big{]}\lesssim\frac{n}{d}\,e^{-\frac{1}{4}d\varepsilon^{3/2}+\varepsilon}\leq\frac{en}{d}\,e^{-\frac{1}{4}d\varepsilon^{3/2}}.$
On the other hand, if $\varepsilon<16d^{-2/3}$ we can estimate
$\mathbf{P}\big{[}\|X\|>2\sqrt{d}\,(1+\varepsilon)\big{]}\leq 1\lesssim\frac{n}{d}\,e^{-\frac{1}{4}d\varepsilon^{3/2}}$
as $\sigma^{2}\leq n\sigma_{*}^{2}=n$ implies $\frac{n}{d}\geq 1$.
Combining the above bounds and using $\sigma^{2}\leq d\leq\sigma^{2}+1$ yields
$\mathbf{P}\big{[}\|X\|>2\sqrt{\sigma^{2}+1}\,(1+\varepsilon)\big{]}\lesssim\frac{n}{\sigma^{2}}\,e^{-\frac{1}{4}\sigma^{2}\varepsilon^{3/2}}$
for $0\leq\varepsilon\leq 1$. This concludes the proof when $\sigma_{*}=1$. For general $\sigma_{*}$, it suffices to apply the above bound to the random matrix $\frac{X}{\sigma_{*}}$. ∎
The rest of the proof of Theorem 3.4 is now identical to that of Theorem 2.3.
3.3. Large deviations
To obtain a large deviations estimate, we proceed as in section 2.3. To this end, we require an analogue of Lemma 2.11 for GOE matrices.
Lemma 3.17.
Let $Y$ be as in Theorem 3.3. For all $t\geq 0$
$\mathbf{E}[\mathop{\mathrm{tr}}e^{tY}]\leq e^{2\sqrt{d}\,t+t^{2}}.$
Proof.
We clearly have $\mathop{\mathrm{tr}}e^{tY}\leq e^{t\lambda_{1}(Y)}$, where $\lambda_{1}(Y)$ denotes the largest eigenvalue of $Y$. It is shown in [8, Theorem 2.11] that $\mathbf{E}[\lambda_{1}(Y)]\leq 2\sqrt{d}$. On the other hand, the Gaussian log-Sobolev inequality implies by [13, eq. (5.8)] and [8, §2.2] that $\lambda_{1}(Y)$ is a $\sqrt{2}$-subgaussian random variable, that is, that
$\mathbf{E}\big{[}e^{t\{\lambda_{1}(Y)-\mathbf{E}[\lambda_{1}(Y)]\}}\big{]}\leq e^{t^{2}}$
for all $t$. Combining these facts yields the conclusion. ∎
The rest of the proof of Theorem 3.5 is now identical to that of Theorem 2.4.
4. The independent case
The aim of this section is to prove the results for the independent entry Model 1.3 that were formulated in the introduction. The main result of this section is the following, where we recall that $\sigma_{1},\sigma_{2},\sigma_{*}$ were defined in (1.2).
Theorem 4.1 (Extremum principle).
Define $X$ as in Model 1.3. Assume that $\sigma_{*}^{2}=1$, and that $\sigma_{1}^{2}\leq d_{1}\in\mathbb{N}$ and $\sigma_{2}^{2}\leq d_{2}\in\mathbb{N}$. Then
$\mathbf{E}[\mathop{\mathrm{tr}}(XX^{*})^{p}]\leq\mathbf{E}[\mathop{\mathrm{tr}}(YY^{*})^{p}]$
for all $p\in\mathbb{N}$, where $Y$ is the $d_{1}\times d_{2}$ matrix with independent entries $Y_{ij}\sim N(0,1)$.
Theorem 1.7 and Corollary 1.8 follow immediately from Theorem 4.1. To deduce Corollary 1.9, note that the block-diagonal matrix $\tilde{Y}$ as illustrated in Figure 1.1 has $\frac{n}{d_{1}}$ blocks, each of which is an independent copy of $Y$. Therefore
$\mathbf{E}[\mathop{\mathrm{Tr}}(XX^{*})^{p}]\leq\frac{n}{d_{1}}\,\mathbf{E}[\mathop{\mathrm{Tr}}(YY^{*})^{p}]=\mathbf{E}[\mathop{\mathrm{Tr}}(\tilde{Y}\tilde{Y}^{*})^{p}]$
is merely another reformulation of Theorem 4.1.
Theorem 4.1 is proved in section 4.1 below, while Theorems 1.4 and 1.6 will be proved in sections 4.2 and 4.3, respectively.
4.1. Extremum principle
Let $(g_{ij})$ be independent $N(0,1)$ random variables. By repeating the proof of Lemma 3.7 verbatim, we may assume without loss of generality in the remainder of this section that $X$ is defined as in Model 1.3 with $\xi_{ij}=g_{ij}$. Defining $E_{ij}:=b_{ij}e_{i}e_{j}^{*}$ for $i\leq n$, $j\leq m$, we can then write
$X=\sum_{i,j}g_{ij}E_{ij}.$
The moments of $X$ are computed as follows.
Lemma 4.2 (Wick formula).
For any $p\in\mathbb{N}$, we have
$\mathbf{E}[\mathop{\mathrm{tr}}(XX^{*})^{p}]=\sum_{\pi\in\mathrm{P}_{2}([2p])}\mathop{\mathrm{tr}}E(\pi)$
with
$E(\pi):=\sum_{(\bm{i},\bm{j})\sim\pi}E_{i_{1}j_{1}}E_{i_{2}j_{2}}^{*}\cdots E_{i_{2p-1}j_{2p-1}}E_{i_{2p}j_{2p}}^{*}.$
Here $\bm{i}\in[n]^{2p},\bm{j}\in[m]^{2p}$, and $(\bm{i},\bm{j})\sim\pi$ denotes $i_{k}=i_{l},j_{k}=j_{l}$ for $\{k,l\}\in\pi$.
Proof.
The proof is identical to that of Lemma 2.5. ∎
The main steps of the proof in this setting are similar to those in the real symmetric case, which we follow with the appropriate modifications. We begin by upper bounding the trace moments by the maximal diagonal entry.
Lemma 4.3.
For any $k\in\mathbb{N}$ and $\pi\in\mathrm{P}_{2}([2k])$, define
$D(\pi):=\max_{r}(E(\pi))_{rr}.$
Then
$\mathbf{E}[\mathop{\mathrm{tr}}(XX^{*})^{p}]\leq\sum_{\pi\in\mathrm{P}_{2}([2p])}D(\pi).$
Moreover, when $b_{ij}=1$ for all $i,j$, the matrix $E(\pi)$ is a multiple of the identity matrix for every $\pi$ and thus the above inequality holds with equality.
Proof.
The proof is identical to that of Lemma 3.9, where we use that when $b_{ij}=1$ the distributions of $X$ and $OX$ coincide for every $O\in O(n)$. ∎
We first consider noncrossing pairings.
Lemma 4.4.
For any $p\in\mathbb{N}$ and noncrossing pairing $\pi\in\mathrm{NC}_{2}([2p])$, we have
$D(\pi)\leq\sigma_{1}^{2\ell(\pi)}\sigma_{2}^{2(p-\ell(\pi))}$
where
$\ell(\pi):=|\{\{i,j\}\in\pi:i\wedge j\text{ is even},~{}i\vee j\text{ is odd}\}|.$
Moreover, equality holds when $b_{ij}=1$ for all $i,j$.
Proof.
As in the proof of Lemma 2.6, there exists $\{k,k+1\}\in\pi$. If $k$ is even, then
$D(\pi)=\max_{r}\sum_{(\bm{i},\bm{j})\sim\pi\backslash\{\{k,k+1\}\}}(E_{i_{1}j_{1}}E_{i_{2}j_{2}}^{*}\cdots E_{i_{k-1}j_{k-1}}\Sigma E_{i_{k+2}j_{k+2}}^{*}\cdots E_{i_{2k}j_{2k}}^{*})_{rr},$
where
$\Sigma=\sum_{i,j}E_{ij}^{*}E_{ij}=\sum_{j}\bigg{(}\sum_{i}b_{ij}^{2}\bigg{)}e_{j}e_{j}^{*}.$
As all $E_{ij}$ have nonnegative entries, we readily estimate
$D(\pi)\leq\sigma_{1}^{2}\,D(\pi\backslash\{\{k,k+1\}\})$
with equality if $b_{ij}=1$ for all $i,j$.
On the other hand, if $k$ is odd, we can repeat the same procedure with
$\Sigma=\sum_{i,j}E_{ij}E_{ij}^{*}=\sum_{i}\bigg{(}\sum_{j}b_{ij}^{2}\bigg{)}e_{i}e_{i}^{*},$
and we obtain
$D(\pi)\leq\sigma_{2}^{2}\,D(\pi\backslash\{\{k,k+1\}\})$
with equality if $b_{ij}=1$ for all $i,j$.
As $\pi\backslash\{\{k,k+1\}\}$ is again a noncrossing pairing, and as removing a consecutive pair from $\pi$ doesn’t change the parity of the indices of the remaining pairs, we can iterate the above argument to conclude the proof. ∎
Before we can analyze the effect of a crossing, we must first obtain the appropriate crossing identities. Let us emphasize that while a noncrossing pairing can only pair even indices with odd indices, a pairing that contains crossings can also pair even indices with each other and odd indices with each other. The appropriate crossing identity depends on the nature of the pairs in the crossing. Here, $M\mathbin{\leq_{\mathrm{e}}}N$ denotes entrywise inequality of matrices, i.e., $M_{ij}\leq N_{ij}$ for all $i,j$.
Lemma 4.5 (Crossing identities and inequalities).
The following hold.
  1. a.
    For any matrices $M_{1},M_{2},M_{3}$ of the appropriate dimensions
    $\sum_{i,j,k,l}E_{ij}M_{1}E_{kl}M_{2}E_{ij}M_{3}E_{kl}=\sum_{i,j,k,l}b_{ij}^{2}b_{kl}^{2}\,(M_{1})_{jk}(M_{2})_{li}(M_{3})_{jk}\,e_{i}e_{l}^{*},$
    $\sum_{i,j,k,l}E_{ij}M_{1}E_{kl}^{*}M_{2}E_{ij}M_{3}E_{kl}^{*}=\sum_{i,j,k,l}b_{ij}^{2}b_{kl}^{2}\,(M_{1})_{jl}(M_{2})_{ki}(M_{3})_{jl}\,e_{i}e_{k}^{*}.$
  2. b.
    For any $m\times n$ matrix $M$
    $\sum_{i,j}\mathop{\mathrm{Tr}}[ME_{ij}]E_{ij}=\sum_{i,j}b_{ij}^{2}\,M_{ji}\,e_{i}e_{j}^{*}.$
  3. c.
    Let $\varepsilon_{1},\ldots,\varepsilon_{4}\in\{1,*\}$. Then for any nonnegative matrices $M_{1},M_{2},M_{3}\mathbin{\geq_{\mathrm{e}}}0$ of the appropriate dimensions, we have
    $\sum_{i,j,k,l}E_{ij}^{\varepsilon_{1}}M_{1}E_{kl}^{\varepsilon_{2}}M_{2}E_{ij}^{\varepsilon_{3}}M_{3}E_{kl}^{\varepsilon_{4}}\mathbin{\leq_{\mathrm{e}}}\sigma_{*}^{4}\begin{cases}M_{2}^{*}M_{3}^{*}M_{1}&\text{ if }\varepsilon_{1}=\varepsilon_{3},\varepsilon_{2}\neq\varepsilon_{4},\\ M_{3}M_{1}^{*}M_{2}^{*}&\text{ if }\varepsilon_{1}\neq\varepsilon_{3},\varepsilon_{2}=\varepsilon_{4},\\ M_{3}M_{2}M_{1}&\text{ if }\varepsilon_{1}\neq\varepsilon_{3},\varepsilon_{2}\neq\varepsilon_{4},\end{cases}$
    with equality if $b_{ij}=1$ for all $i,j$.
Proof.
All parts follow immediately from the definition of $E_{ij}$. ∎
We now introduce the relevant crossing types in this setting.
Definition 4.6.
Let $\pi\in\mathrm{P}_{2}([2p])$, and fix a crossing $\{i,k\},\{j,l\}\in\pi$ such that $i<j<k<l$. Then this crossing is said to be of
  1. $\bullet$
    type 1 if $i,k$ have opposite parity or $j,l$ have opposite parity.
  2. $\bullet$
    type 2 if $i,k$ have the same parity, $j,l$ have the same parity, and there exists a pair $\{a,b\}\in\pi$ with $a\in(i,j)\cup(k,l)$ and $b\not\in(i,j)\cup(k,l)$;
  3. $\bullet$
    type 3 otherwise.
We now consider each crossing type in turn.
Lemma 4.7 (Type 1 crossing).
Let $p\in\mathbb{N}$, $\pi\in\mathrm{P}_{2}([2p])$, and $\{a,c\},\{b,d\}\in\pi$ with $a<b<c<d$ be a crossing of type 1. Then there exists a pairing $\pi^{\prime}\in\mathrm{P}_{2}([2p-4])$, whose definition depends only on $\pi,a,b,c,d$, so that
$D(\pi)\leq\sigma_{*}^{4}D(\pi^{\prime}).$
Moreover, equality holds when $b_{ij}=1$ for all $i,j$.
Proof.
For simplicity, consider the case that $a,b,c$ are odd and $d$ is even; the proof in the remaining cases is identical. Then Lemma 4.5(c) allows us to estimate
$D(\pi)$ $=\max_{r}\sum_{(\bm{i},\bm{j})\sim\pi\backslash\{\{a,c\},\{b,d\}\}}\sum_{i,j,k,l}(M_{0}E_{ij}M_{1}E_{kl}M_{2}E_{ij}M_{3}E_{kl}^{*}M_{4})_{rr}$
$\leq\sigma_{*}^{4}\,\max_{r}\sum_{(\bm{i},\bm{j})\sim\pi\backslash\{\{a,c\},\{b,d\}\}}(M_{0}M_{2}^{*}M_{3}^{*}M_{1}M_{4})_{rr}$
with $M_{0}:=E_{i_{1}j_{1}}E_{i_{2}j_{2}}^{*}\cdots E_{i_{a-1}j_{a-1}}^{*}$, $M_{1}:=E_{i_{a+1}j_{a+1}}^{*}E_{i_{a+2}j_{a+2}}\cdots E_{i_{b-1}j_{b-1}}^{*}$, etc. We readily read off the existence of $\pi^{\prime}\in\mathrm{P}_{2}([2p-4])$ so that the right-hand side equals $\sigma_{*}^{4}D(\pi^{\prime})$. Moreover, Lemma 4.5(c) ensures equality when $b_{ij}=1$ for all $i,j$. ∎
Lemma 4.8 (Type 2 crossing).
Let $p\in\mathbb{N}$, $\pi\in\mathrm{P}_{2}([2p])$, $\{a,c\},\{b,d\},\{e,f\}\in\pi$ where $a<b<c<d$, $a,c$ have the same parity, $b,d$ have the same parity, and $e\in(a,b)\cup(c,d)$, $f\not\in(a,b)\cup(c,d)$. Then there exists a pairing $\pi^{\prime}\in\mathrm{P}_{2}([2p-6])$, whose definition depends only on $\pi,a,b,c,d,e,f$, so that
$D(\pi)\leq\sigma_{*}^{6}D(\pi^{\prime}).$
Moreover, equality holds when $b_{ij}=1$ for all $i,j$.
Proof.
Let us first assume that $a,b,c,d$ are all odd. Then the first equation display of Lemma 4.5(a) enables us to estimate
$D(\pi)\leq\sigma_{*}^{4}\,\max_{r}\sum_{(\bm{i},\bm{j})\sim\pi\backslash\{\{a,c\},\{b,d\}\}}\mathop{\mathrm{Tr}}[M_{1}^{*}M_{3}](M_{0}M_{2}^{*}M_{4})_{rr}$
with equality if $b_{ij}=1$ for all $i,j$, where $M_{k}$ are as in the proof of Lemma 4.7. Now note that, by assumption, $\{e,f\}$ pairs a term inside the trace with a term inside the matrix element on the right-hand side. We can therefore use Lemma 4.5(b) or its adjoint (depending on the parities of $e$ and $f$) to estimate the right-hand side by $\sigma_{*}^{6}D(\pi^{\prime})$ for some $\pi^{\prime}\in\mathrm{P}_{2}([2p-6])$ that depends only on $\pi,a,b,c,d,e,f$, with equality if $b_{ij}=1$ for all $i,j$. We omit the details which are completely analogous to the corresponding argument in the proof of Lemma 3.13.
The other possible parities of $a,b,c,d$ are treated in a completely analogous way: if $a,c$ are odd and $b,d$ are even we use the second equation display of Lemma 4.5(a) instead of the first, while the remaining two cases use the adjoint of the first or second equation display of Lemma 4.5(a), respectively. ∎
Lemma 4.9 (Type 3 crossing).
Let $p\in\mathbb{N}$, $\pi\in\mathrm{P}_{2}([2p])$, and let $\{a,c\},\{b,d\}\in\pi$ with $a<b<c<d$ be a crossing of type 3. Then there exist $\pi_{1}\in\mathrm{P}_{2}([b-a+d-c-2])$ and $\pi_{2}\in\mathrm{P}_{2}([2p-2-b+a-d+c])$, which depend only on $\pi,a,b,c,d$, so that
$D(\pi)\leq\begin{cases}\sigma_{1}^{2}\sigma_{*}^{2}D(\pi_{1})D(\pi_{2})&\text{if }d\text{ is odd},\\ \sigma_{2}^{2}\sigma_{*}^{2}D(\pi_{1})D(\pi_{2})&\text{ if }d\text{ is even}.\end{cases}$
Moreover, equality holds when $b_{ij}=1$ for all $i,j$.
Proof.
Let us first assume that $a,b,c,d$ are all odd. Then the first equation display of Lemma 4.5(a) enables us to estimate
$D(\pi)$ $=\max_{r}\sum_{(\bm{i},\bm{j})\sim\pi\backslash\{\{a,c\},\{b,d\}\}}\sum_{i,j,k,l}b_{ij}^{2}b_{kl}^{2}(M_{0})_{ri}(M_{1})_{jk}(M_{2})_{li}(M_{3})_{jk}(M_{4})_{lr}$
$\leq\sigma_{*}^{2}\,\max_{r}\sum_{l}\Bigg{(}\sum_{(\bm{i}_{J},\bm{j}_{J})\sim\pi_{J}}(M_{0}M_{2}^{*})_{rl}(M_{4})_{lr}\Bigg{)}\Bigg{(}\sum_{k}b_{kl}^{2}\sum_{(\bm{i}_{I},\bm{j}_{I})\sim\pi_{I}}(M_{1}^{*}M_{3})_{kk}\Bigg{)}$
with equality when $b_{ij}=1$ for all $i,j$, where $M_{k}$ are as in the proof of Lemma 4.7 and $\pi\backslash\{\{a,c\},\{b,d\}\}=\pi_{I}\cup\pi_{J}$ as in the proof of Lemma 3.14. The conclusion now follows readily by the same argument as in the proof of Lemma 3.14.
The other possible parities of $a,b,c,d$ are treated in a completely analogous way, where we must take care to observe that we gain a factor $\sigma_{1}^{2}$ or $\sigma_{2}^{2}$ depending on whether $b,d$ are odd or even, respectively. ∎
We can now complete the proof of Theorem 4.1.
Proof of Theorem 4.1.
We first estimate $\mathbf{E}[\mathop{\mathrm{tr}}(XX^{*})^{p}]$ as in Lemma 4.3. We then repeatedly apply the following to each quantity $D(\pi)$ in the resulting expression:
  1. $\bullet$
    If $\pi$ contains a crossing, we apply either Lemma 4.7, 4.8, or 4.9 to the smallest crossing in the lexicographic order, depending on whether that crossing is of type 1, 2, or 3, respectively. In the type 2 case, we choose the smallest pair $\{e,f\}$ in the lexicographic order that satisfies the assumption of Lemma 4.8.
  2. $\bullet$
    If $\pi$ is noncrossing, we apply Lemma 4.4.
This algorithm gives rise to an inequality of the form
$\mathbf{E}[\mathop{\mathrm{tr}}(XX^{*})^{p}]\leq\sum_{k,l\geq 0:k+l\leq p}\tilde{\varkappa}_{p}(k,l)\sigma_{1}^{2k}\sigma_{2}^{2l}\sigma_{*}^{2p-2k-2l},$
where the coefficients $\tilde{\varkappa}_{p}(k,l)\in\mathbb{Z}_{+}$ are independent of the matrix $X$.
Moreover, as all the inequalities used above become equalities when $b_{ij}=1$ for all $i,j$, we can apply the same algorithm to compute
$\mathbf{E}[\mathop{\mathrm{tr}}(YY^{*})^{p}]=\sum_{k,l\geq 0:k+l\leq p}\tilde{\varkappa}_{p}(k,l)d_{1}^{k}d_{2}^{l}.$
The conclusion readily follows from the assumptions $\sigma_{*}=1$, $\sigma_{1}^{2}\leq d_{1}$, $\sigma_{2}^{2}\leq d_{2}$. ∎
4.2. Small deviations
The difficulty as compared with the self-adjoint models discussed in the previous sections is to obtain the following moment estimate.
Lemma 4.10.
Define $Y$ as in Theorem 4.1, and assume that $d_{1}\leq d_{2}$. Then
$\mathbf{E}[\mathop{\mathrm{tr}}(YY^{*})^{p}]\lesssim\frac{1}{d_{1}}\big{(}d_{1}^{1/2}+d_{2}^{1/2}\big{)}^{2p}\bigg{(}1+\frac{8p^{2}}{d_{1}^{1/2}d_{2}^{3/2}}\bigg{)}^{p}$
for all $p\geq d_{1}^{1/6}d_{2}^{1/2}$.
Proof.
The result follows immediately from Theorem 5.2 in section 5 below. ∎
We can now essentially repeat the proof of Proposition 3.16.
Proposition 4.11.
For $X$ as in Model 1.3 with $\sigma_{1}\leq\sigma_{2}$, we have
$\mathbf{P}\big{[}\|X\|>\big{(}{\textstyle\sqrt{\sigma_{1}^{2}+\sigma_{*}^{2}}+\sqrt{\sigma_{2}^{2}+\sigma_{*}^{2}}}\big{)}(1+\varepsilon)\big{]}\leq\frac{n\sigma_{*}^{2}}{C\sigma_{1}^{2}}e^{-\frac{1}{8}\frac{\sigma_{1}^{1/2}\sigma_{2}^{3/2}}{\sigma_{*}^{2}}\varepsilon^{3/2}}$
for all $0\leq\varepsilon\leq 1$, where $C$ is a universal constant.
Proof.
Suppose first that $\sigma_{*}=1$, and let $d_{1}=\lceil\sigma_{1}^{2}\rceil$ and $d_{2}=\lceil\sigma_{2}^{2}\rceil$. Using Markov’s inequality, $\|X\|^{2p}\leq n\mathop{\mathrm{tr}}(XX^{*})^{p}$, Theorem 4.1, and Lemma 4.10, we obtain
$\mathbf{P}\big{[}\|X\|\geq(d_{1}^{1/2}+d_{2}^{1/2})(1+\varepsilon)\big{]}\lesssim\frac{n}{d_{1}}e^{-\varepsilon p+\frac{8p^{3}}{d_{1}^{1/2}d_{2}^{3/2}}}$
for all $p\geq d_{1}^{1/6}d_{2}^{1/2}$, where we used $(1+\varepsilon)^{-2p}\leq e^{-\varepsilon p}$ for $0<\varepsilon\leq 1$ and $1+x\leq e^{x}$. If $\varepsilon\geq 64d_{1}^{-1/6}d_{2}^{-1/2}$, we may choose $p=\lfloor\frac{1}{4}d_{1}^{1/4}d_{2}^{3/4}\sqrt{\varepsilon}\rfloor\geq d_{1}^{1/6}d_{2}^{1/2}$ to obtain
$\mathbf{P}\big{[}\|X\|\geq(d_{1}^{1/2}+d_{2}^{1/2})(1+\varepsilon)\big{]}\lesssim\frac{en}{d_{1}}e^{-\frac{1}{8}d_{1}^{1/4}d_{2}^{3/4}\varepsilon^{3/2}}.$
If $\varepsilon<64d_{1}^{-1/6}d_{2}^{-1/2}$, the same bound is valid as
$\mathbf{P}\big{[}\|X\|\geq(d_{1}^{1/2}+d_{2}^{1/2})(1+\varepsilon)\big{]}\leq 1\lesssim\frac{n}{d_{1}}e^{-\frac{1}{8}d_{1}^{1/4}d_{2}^{3/4}\varepsilon^{3/2}}$
using that $\sigma_{1}^{2}\leq n\sigma_{*}^{2}$ implies $\frac{n}{d_{1}}\geq 1$.
Combining the above bounds readily yields
$\mathbf{P}\big{[}\|X\|\geq\big{(}{\textstyle\sqrt{\sigma_{1}^{2}+1}+\sqrt{\sigma_{2}^{2}+1}}\big{)}(1+\varepsilon)\big{]}\lesssim\frac{n}{\sigma_{1}^{2}}e^{-\frac{1}{8}\sigma_{1}^{1/2}\sigma_{2}^{3/2}\varepsilon^{3/2}}$
for $0\leq\varepsilon\leq 1$. This concludes the proof when $\sigma_{*}=1$. For general $\sigma_{*}$, it suffices to apply the above bound to the random matrix $\frac{X}{\sigma_{*}}$. ∎
We now complete the proof of Theorem 1.4.
Proof of Theorem 1.4.
Proposition 4.11 yields
$\mathbf{P}\big{[}\|X\|>\sigma_{1}+\sigma_{2}+\tfrac{\sigma_{*}^{2}}{\sigma_{1}}+\tfrac{3}{4}\sigma_{*}^{4/3}\sigma_{1}^{-1/3}t\big{]}\leq\frac{n\sigma_{*}^{2}}{C\sigma_{1}^{2}}e^{-\frac{1}{64}t^{3/2}}$
by setting $\varepsilon=\frac{1}{4}t\sigma_{*}^{4/3}\sigma_{1}^{-1/3}\sigma_{2}^{-1}$, where we used that $\sigma_{1}\leq\sigma_{2}$ so that
${\textstyle\sqrt{\sigma_{1}^{2}+\sigma_{*}^{2}}+\sqrt{\sigma_{2}^{2}+\sigma_{*}^{2}}}\leq\sigma_{1}+\sigma_{2}+\frac{\sigma_{*}^{2}}{\sigma_{1}}\leq 3\sigma_{2}.$
The conclusion is immediate for $t\geq\tfrac{4\sigma_{*}^{2/3}}{\sigma_{1}^{2/3}}$, and follows for $t<\tfrac{4\sigma_{*}^{2/3}}{\sigma_{1}^{2/3}}\leq 4$ as then
$1\leq\frac{n\sigma_{*}^{2}}{\sigma_{1}^{2}}\leq\frac{en\sigma_{*}^{2}}{\sigma_{1}^{2}}e^{-\frac{1}{64}t^{3/2}}.$
This concludes the proof. ∎
4.3. Large deviations
We require the following analogue of Lemma 3.17.
Lemma 4.12.
Let $Y$ be as in Theorem 4.1. For all $t\geq 0$
$\mathbf{E}\big{[}\mathop{\mathrm{tr}}e^{t(YY^{*})^{1/2}}\big{]}\leq e^{(\sqrt{d_{1}}+\sqrt{d_{2}})t+t^{2}/2}.$
Proof.
We clearly have $\mathop{\mathrm{tr}}e^{t(YY^{*})^{1/2}}\leq e^{t\|Y\|}$. It is shown in [8, Theorem 2.13] that $\mathbf{E}\|Y\|\leq\sqrt{d_{1}}+\sqrt{d_{2}}$. The Gaussian log-Sobolev inequality implies by [13, eq. (5.8)] and [8, §2.2] that $\|Y\|$ is a $1$-subgaussian random variable, that is, that
$\mathbf{E}\big{[}e^{t\{\|Y\|-\mathbf{E}\|Y\|\}}\big{]}\leq e^{t^{2}/2}$
for all $t$. Combining these facts yields the conclusion. ∎
This yields the following.
Proposition 4.13.
For $X$ as in Model 1.3 with $n\leq m$, we have
$\mathbf{P}\big{[}\|X\|>{\textstyle\sqrt{\sigma_{1}^{2}+\sigma_{*}^{2}}+\sqrt{\sigma_{2}^{2}+\sigma_{*}^{2}}}+\sigma_{*}\varepsilon\big{]}\leq 2n\,e^{-\varepsilon^{2}/2}$
for all $\varepsilon\geq 0$.
Proof.
Suppose first that $\sigma_{*}=1$, and let $d_{1}=\lceil\sigma_{1}^{2}\rceil$ and $d_{2}=\lceil\sigma_{2}^{2}\rceil$. Then
$\frac{1}{2}\,\mathbf{E}\big{[}\mathop{\mathrm{tr}}e^{t(XX^{*})^{1/2}}\big{]}$ $\leq\mathbf{E}[\mathop{\mathrm{tr}}\cosh(t(XX^{*})^{1/2})]$
$\leq\mathbf{E}[\mathop{\mathrm{tr}}\cosh(t(YY^{*})^{1/2})]\leq\mathbf{E}\big{[}\mathop{\mathrm{tr}}e^{t(YY^{*})^{1/2}}\big{]},$
by Theorem 4.1, where we used that the Taylor expansion of the hyperbolic cosine only has terms of even degree. As $e^{t\|X\|}\leq n\mathop{\mathrm{tr}}e^{t(XX^{*})^{1/2}}$, we obtain
$\mathbf{E}[e^{t\|X\|}]\leq 2n\,e^{(\sqrt{d_{1}}+\sqrt{d_{2}})t+t^{2}/2}$
using Lemma 4.12. By Markov’s inequality
$\mathbf{P}\big{[}\|X\|>\sqrt{d_{1}}+\sqrt{d_{2}}+\varepsilon\big{]}\leq\frac{\mathbf{E}[e^{t\|X\|}]}{e^{(\sqrt{d_{1}}+\sqrt{d_{2}}+\varepsilon)t}}\leq 2n\,e^{-\varepsilon t+t^{2}/2}.$
Optimizing over $t\geq 0$ yields
$\mathbf{P}\big{[}\|X\|>{\textstyle\sqrt{\sigma_{1}^{2}+1}+\sqrt{\sigma_{2}^{2}+1}}+\varepsilon\big{]}\leq 2n\,e^{-\varepsilon^{2}/2}$
for all $\varepsilon\geq 0$. This concludes the proof for $\sigma_{*}=1$. For general $\sigma_{*}$, it suffices to apply the above bound to the random matrix $\frac{X}{\sigma_{*}}$. ∎
Theorem 1.6 follows readily.
Proof of Theorem 1.6.
The conclusion follows immediately from Proposition 4.13 using that $\sqrt{\sigma_{1}^{2}+\sigma_{*}^{2}}+\sqrt{\sigma_{2}^{2}+\sigma_{*}^{2}}\leq\sigma_{1}+\sigma_{2}+\frac{\sigma_{*}^{2}}{2\sigma_{1}}+\frac{\sigma_{*}^{2}}{2\sigma_{2}}\leq\sigma_{1}+\sigma_{2}+\sigma_{*}$. ∎
5. Moment estimates for Wishart matrices
The aim of this section is to prove the moment estimate for Wishart matrices that was used in Lemma 4.10 above. Throughout this section, we fix
$n\leq m,\qquad\qquad c:=\frac{m}{n}\geq 1.$
We let $Y$ be an $n\times m$ matrix with i.i.d. $N_{\mathbb{R}}(0,1)$ entries, and $Z$ be an $n\times m$ matrix with i.i.d. $N_{\mathbb{C}}(0,1)$ entries. Our main results are as follows.
Theorem 5.1 (Complex Wishart moments).
For $p\in\mathbb{N}$, we have
$\mathbf{E}[\mathop{\mathrm{tr}}(ZZ^{*})^{p}]\lesssim n^{p}(\sqrt{c}+1)^{2p}\bigg{(}1+\frac{2p^{2}}{c^{3/2}n^{2}}\bigg{)}^{p}\frac{c^{3/4}}{p^{3/2}}.$
Theorem 5.2 (Real Wishart moments).
For $p\in\mathbb{N}$, we have
$\mathbf{E}[\mathop{\mathrm{tr}}(YY^{*})^{p}]\lesssim n^{p}(\sqrt{c}+1)^{2p}\bigg{(}1+\frac{8p^{2}}{c^{3/2}n^{2}}\bigg{)}^{p}\bigg{(}\frac{c^{3/4}}{p^{3/2}}+\frac{1}{n}\bigg{)}.$
A moment estimate for complex Wishart matrices was previously obtained by Ledoux in [14, p. 201]. However, the constants in Ledoux’ estimate depend in an unspecified manner on the aspect ratio $c$, making it unsuitable for the purposes of this paper. [^3] The argument of [14, p. 201] also contains a further issue, that the recursion in eq. (29) of that paper does not imply the inequality for $b_{p}$ claimed subsequently. The difficulty in the proofs of Theorems 5.1 and 5.2 is to obtain bounds that have optimal dependence on $c$, which requires a more delicate understanding of the structure of the associated moment recursions.
Remark 5.3.
To illustrate the sharpness of Theorems 5.1 and 5.2, let us make two observations. First, the argument of section 4.2 shows that these moment estimates yield tail bounds as in Theorem 1.4, which match the exact Tracy-Widom asymptotics (1.1) with the optimal order of the fluctuations and tail behavior.
On the other hand, if we let $n\to\infty$ with $p,c$ fixed, it is classical [18, p. 368] that both $n^{-p}\mathbf{E}[\mathop{\mathrm{tr}}(YY^{*})^{p}]$ and $n^{-p}\mathbf{E}[\mathop{\mathrm{tr}}(ZZ^{*})^{p}]$ converge to the $p$-moment $\chi_{p}^{c}$ of the Marchenko–Pastur distribution. By using the explicit formula for its generating function [18, p. 205], the Darboux method [23, Theorem 5.11] yields
$\chi_{p}^{c}=(1+o(1))\,(\sqrt{c}+1)^{2p}\,\frac{c^{1/4}(\sqrt{c}+1)}{2\sqrt{\pi}\,p^{3/2}}$
as $p\to\infty$. The estimates of Theorems 5.1 and 5.2 reproduce the exact asymptotics of $\chi_{p}^{c}$ precisely up to a universal constant.
The remainder of this section is organized as follows. In section 5.1, we recall the recursive formulas of [10, 7] for the moments of complex and real Wishart matrices. Let us emphasize at the outset that the recursive formula for real Wishart moments involves complex Wishart moments, so that we must consider the latter even if one is ultimately only interested in real Wishart matrices. We then prove Theorem 5.1 in section 5.2, and finally prove Theorem 5.2 in section 5.3.
5.1. Moment recursions
In addition to the random matrices $Y$ and $Z$ defined above, we also introduce an $(n-1)\times(m-1)$ matrix $Z^{\prime}$ with i.i.d. $N_{\mathbb{C}}(0,1)$ entries. Throughout the remainder of this section, we define
$A_{p}:=\frac{\mathbf{E}[\mathop{\mathrm{Tr}}(ZZ^{*})^{p}]}{n^{p+1}},\qquad A_{p}^{\prime}:=\frac{\mathbf{E}[\mathop{\mathrm{Tr}}(Z^{\prime}Z^{\prime*})^{p}]}{n^{p+1}},\qquad B_{p}:=\frac{\mathbf{E}[\mathop{\mathrm{Tr}}(YY^{*})^{p}]}{n^{p+1}},$
so that $\mathbf{E}[\mathop{\mathrm{tr}}(ZZ^{*})^{p}]=n^{p}A_{p}$ and $\mathbf{E}[\mathop{\mathrm{tr}}(YY^{*})^{p}]=n^{p}B_{p}$.
The proofs of the main results of this section are based on recursive formulas for $A_{p},A_{p}^{\prime},B_{p}$ that we presently recall. The following recursive formula for $A_{p}$ was obtained by Haagerup and Thorbjørnsen [10, Theorem 8.2].
Theorem 5.4 (Haagerup-Thorbjørnsen).
For $p\geq 1$
$A_{p+1}=2(c+1)\,\frac{2p+1}{2p+4}\,A_{p}-\bigg{(}(c-1)^{2}-\frac{p^{2}}{n^{2}}\bigg{)}\,\frac{p-1}{p+2}\,A_{p-1},$
with the initial conditions $A_{0}=1$ and $A_{1}=c$.
A direct consequence is the following.
Corollary 5.5.
For $p\geq 1$
$A_{p+1}^{\prime}=2\bigg{(}c+1-\frac{2}{n}\bigg{)}\,\frac{2p+1}{2p+4}\,A_{p}^{\prime}-\bigg{(}(c-1)^{2}-\frac{p^{2}}{n^{2}}\bigg{)}\,\frac{p-1}{p+2}\,A_{p-1}^{\prime},$
with $A_{0}^{\prime}=\frac{n-1}{n}$ and $A_{1}^{\prime}=\frac{n-1}{n}(c-\frac{1}{n})$. Moreover, $A_{p}^{\prime}\leq A_{p}$ for all $p$.
Proof.
The recursion for $A_{p}^{\prime}$ follows readily from Theorem 5.4. The inequality $A_{p}^{\prime}\leq A_{p}$ follows from Jensen’s inequality by the convexity of $Z\mapsto\mathop{\mathrm{Tr}}(ZZ^{*})^{p}$ by conditioning on the $(n-1)\times(m-1)$ principal submatrix of $Z$. ∎
The reason that we consider the modified moments $A_{p}^{\prime}$ is that they appear in the following moment recursion for $B_{p}$ due to Cunden et al. [7, Theorem 3.5] (the values of $B_{1},B_{2}$ stated here are readily obtained by a straightforward computation).
Theorem 5.6 (Cunden et al.).
For $p\geq 2$
$B_{p+1}=2\bigg{(}c+1-\frac{1}{n}\bigg{)}B_{p}-\bigg{(}(c-1)^{2}-\frac{4p(p-1)+1}{n^{2}}\bigg{)}B_{p-1}\\ +\frac{3}{p-1}\bigg{[}\bigg{(}c+1-\frac{p+1}{n}\bigg{)}A_{p}^{\prime}-A_{p+1}^{\prime}\bigg{]},$
with $B_{0}=1$, $B_{1}=c$, and $B_{2}=(c+1+\frac{1}{n})c$.
Finally, we will define in the sequel $\chi_{p}:=4^{-p}C_{p}$, where $C_{p}$ denotes the $p$th Catalan number. We recall the well known Catalan recursion
$\chi_{p+1}=\frac{2p+1}{2p+4}\,\chi_{p},\qquad\quad\chi_{0}=1$ (5.1)
for $p\geq 0$, as well as the standard estimate $\chi_{p}\lesssim p^{-3/2}$ by Stirling’s formula.
5.2. Complex Wishart moments
The aim of this section is to bound the complex Wishart moments $A_{p}$. To this end, it will be convenient to define
$K_{p}:=\frac{A_{p+1}}{A_{p}}\frac{\chi_{p}}{\chi_{p+1}},$
so that
$A_{p}=4c\,K_{1}K_{2}\cdots K_{p-1}\chi_{p}.$
It follows readily from Theorem 5.4 and (5.1) that
$K_{p}=2(c+1)-\bigg{(}1-\frac{3}{4p^{2}-1}\bigg{)}\bigg{(}(c-1)^{2}-\frac{p^{2}}{n^{2}}\bigg{)}\frac{1}{K_{p-1}}$ (5.2)
for $p\geq 1$. Note that this recursion does not require an initial condition, as the second term on the right-hand side vanishes for $p=1$.
The analysis of this equation depends on the sign of the second term on the right-hand side. We consider the two cases $p<(c-1)n$ and $p\geq(c-1)n$ separately.
5.2.1. The case $p<(c-1)n$
In this subsection, we fix $1\leq p<(c-1)n$ and let
$\lambda:=c+1+\sqrt{4c+\frac{p^{2}}{n^{2}}},\qquad\qquad\bar{\lambda}:=c+1-\sqrt{4c+\frac{p^{2}}{n^{2}}}.$
Note that $2(c+1)=\lambda+\bar{\lambda}$ and $(c-1)^{2}-\frac{p^{2}}{n^{2}}=\lambda\bar{\lambda}$. In particular, the latter implies that $\bar{\lambda}>0$ as we assumed that $p<(c-1)n$. We can therefore estimate
$K_{k}\leq\lambda+\bar{\lambda}-\bigg{(}1-\frac{3}{4k^{2}-1}\bigg{)}\frac{\lambda\bar{\lambda}}{K_{k-1}}$ (5.3)
for all $1\leq k\leq p$ using (5.2).
At the core of the proof lie two distinct bounds on $K_{k}$ for $1\leq k\leq p$. The first bound will be used for large $k$, while the second bound will be used for small $k$.
Lemma 5.7.
For $1\leq k\leq p$, we have
$K_{k}\leq\bigg{(}1+\frac{2\sqrt{c}}{k^{2}}\bigg{)}\lambda.$
Proof.
First, note that $K_{1}=2(c+1)\leq 2\lambda\leq(1+2\sqrt{c})\lambda$ as $c\geq 1$. For $k>1$, we proceed by induction. Assuming we have proved the result for $k\leftarrow k-1$, we have
$K_{k}\leq\lambda+\bar{\lambda}-\frac{1-\frac{3}{4k^{2}-1}}{1+\frac{2\sqrt{c}}{(k-1)^{2}}}\bar{\lambda}.$
using (5.3). To conclude the result, we must therefore show that
$\frac{2\sqrt{c}}{(k-1)^{2}}+\frac{3}{4k^{2}-1}\leq\frac{2\sqrt{c}}{k^{2}}\bigg{(}1+\frac{2\sqrt{c}}{(k-1)^{2}}\bigg{)}\frac{\lambda}{\bar{\lambda}}.$
Subtracting $\frac{2\sqrt{c}}{k^{2}}$ on both sides and rearranging yields
$2\sqrt{c}\,(2k-1)+\frac{3k^{2}(k-1)^{2}}{4k^{2}-1}\leq 2\sqrt{c}\,(k-1)^{2}\,\frac{\lambda-\bar{\lambda}}{\bar{\lambda}}+4c\,\frac{\lambda}{\bar{\lambda}}.$
But as
$\frac{\lambda-\bar{\lambda}}{\bar{\lambda}}\geq\frac{4}{\sqrt{c}},\qquad\quad\frac{\lambda}{\bar{\lambda}}\geq 1,\qquad\quad 4k^{2}-1\geq 4k(k-1),$
it suffices to prove the quadratic inequality for $k-1$
$7(k-1)^{2}-\bigg{(}4\sqrt{c}+\frac{3}{4}\bigg{)}(k-1)+4c-2\sqrt{c}\geq 0$
Thus it suffices to check that the quadratic function has nonpositive discriminant, which is readily verified using that $c\geq 1$. ∎
Lemma 5.8.
For $1\leq k\leq p$, we have
$K_{k}\leq\bigg{(}1+\frac{3}{2k}\bigg{)}\lambda.$
Proof.
Clearly $K_{1}\leq 2\lambda\leq(1+\frac{3}{2})\lambda$. For $k>1$ we proceed again by induction. Assuming we have proved the result for $k\leftarrow k-1$, we have
$K_{k}\leq\lambda+\bar{\lambda}-\frac{1-\frac{3}{4k^{2}-1}}{1+\frac{3}{2(k-1)}}\bar{\lambda}$
using (5.3). We must therefore show that
$\frac{3}{2(k-1)}+\frac{3}{4k^{2}-1}\leq\frac{3}{2k}\bigg{(}1+\frac{3}{2(k-1)}\bigg{)}\frac{\lambda}{\bar{\lambda}}.$
Using that $\lambda\geq\bar{\lambda}$ and rearranging, it suffices to show that
$\frac{1}{2k(k-1)}+\frac{1}{4k^{2}-1}\leq\frac{3}{4k(k-1)}.$
This is always true as $4k^{2}-1\geq 4k(k-1)$. ∎
Combining the above bounds yields the following conclusion.
Proposition 5.9.
For $1\leq p<(c-1)n$, we have
$A_{p}\lesssim(\sqrt{c}+1)^{2p}\bigg{(}1+\frac{p^{2}}{4c^{3/2}n^{2}}\bigg{)}^{p}\frac{c^{3/4}}{p^{3/2}}.$
Proof.
Note that Lemma 5.8 implies
$K_{1}\cdots K_{\lfloor\sqrt{c}\rfloor}\leq\lambda^{\lfloor\sqrt{c}\rfloor}\prod_{k=1}^{\lfloor\sqrt{c}\rfloor}\bigg{(}1+\frac{3}{2k}\bigg{)}\lesssim c^{3/4}\lambda^{\lfloor\sqrt{c}\rfloor},$
while Lemma 5.7 implies
$K_{\lfloor\sqrt{c}\rfloor+1}\cdots K_{p-1}\leq\lambda^{p-1-\lfloor\sqrt{c}\rfloor}\prod_{k=\lfloor\sqrt{c}\rfloor+1}^{\infty}\bigg{(}1+\frac{2\sqrt{c}}{k^{2}}\bigg{)}\lesssim\lambda^{p-1-\lfloor\sqrt{c}\rfloor}.$
We therefore have
$A_{p}=4c\,K_{1}\cdots K_{p-1}\chi_{p}\lesssim c^{3/4}\lambda^{p}\chi_{p},$
where we used $\lambda\geq c$. It remains to note that
$\lambda-(\sqrt{c}+1)^{2}=\sqrt{4c+\frac{p^{2}}{n^{2}}}-\sqrt{4c}\leq\frac{1}{4\sqrt{c}}\frac{p^{2}}{n^{2}}\leq\frac{p^{2}}{4c^{3/2}n^{2}}(\sqrt{c}+1)^{2}$
and that $\chi_{p}\lesssim p^{-3/2}$. ∎
Remark 5.10.
It is instructive to note the features of the analysis that were needed to obtain a sharp bound. In Lemma 5.7, the constant $2$ is unimportant but the correct dependence on $c$ is key. In contrast, the optimal constant $\frac{3}{2}$ in Lemma 5.8 is used in a crucial way to obtain the correct exponent $c^{3/4}$ in the final bound.
5.2.2. The case $p\geq(c-1)n$
This case is much easier and yields a qualitatively better bound (the latter is however irrelevant for our purposes).
Proposition 5.11.
For $p\geq(c-1)n$, we have
$A_{p}\lesssim(\sqrt{c}+1)^{2p}\bigg{(}1+\frac{2p^{2}}{c^{2}n^{2}}\bigg{)}^{p}\frac{1}{p^{3/2}}.$
Proof.
The assumption $p\geq(c-1)n$ implies
$2(c+1)=(\sqrt{c}+1)^{2}+\frac{(c-1)^{2}}{(\sqrt{c}+1)^{2}}\leq(\sqrt{c}+1)^{2}+\frac{p^{2}}{cn^{2}}.$
Define $N_{k}:=\frac{A_{k}}{\chi_{k}}$. Then we can crudely estimate for $1\leq k\leq p$
$N_{k+1}\leq(\sqrt{c}+1)^{2}N_{k}+\frac{p^{2}}{cn^{2}}\,N_{k}+1_{k>1}\,\frac{p^{2}}{n^{2}}\,N_{k-1}$
using Theorem 5.4 and (5.1). To conclude the proof, it suffices to show this implies
$N_{k}\leq 4(\sqrt{c}+1)^{2k}\bigg{(}1+\frac{2p^{2}}{c^{2}n^{2}}\bigg{)}^{k}$
for all $1\leq k\leq p$. The claim is trivial for $k=1$ as $N_{1}=4c$, while the claim is readily verified to hold for $k>1$ by induction. ∎
The proof of Theorem 5.1 is now immediate.
Proof of Theorem 5.1.
Combining Propositions 5.9 and 5.11 yields
$A_{p}\lesssim(\sqrt{c}+1)^{2p}\bigg{(}1+\frac{2p^{2}}{c^{3/2}n^{2}}\bigg{)}^{p}\frac{c^{3/4}}{p^{3/2}}$
using $c\geq 1$. The conclusion follows from the definition of $A_{p}$. ∎
5.3. Real Wishart moments
Taking inspiration from [16], we define
$D_{p}:=B_{p}-A_{p}^{\prime}.$
Then we have the following.
Lemma 5.12.
For all $p\geq 1$, we have $D_{p}\geq 0$ and
$D_{p+1}=2\bigg{(}c+1-\frac{1}{n}\bigg{)}D_{p}-\bigg{(}(c-1)^{2}-\frac{4p(p-1)+1}{n^{2}}\bigg{)}D_{p-1}\\ -\frac{1}{n}A_{p}^{\prime}+\frac{(3p-1)(p-1)}{n^{2}}A_{p-1}^{\prime},$
with the initial conditions $D_{0}=\frac{1}{n}$ and $D_{1}=\frac{1}{n}(c+1-\frac{1}{n})$.
Proof.
To show $D_{p}\geq 0$, it suffices by Corollary 5.5 to show that $D_{p}\geq B_{p}-A_{p}\geq 0$, that is, that $\mathbf{E}[\mathop{\mathrm{Tr}}(ZZ^{*})^{p}]\leq\mathbf{E}[\mathop{\mathrm{Tr}}(YY^{*})^{p}]$. This follows from the Wick formula, as $\mathbf{E}[\mathop{\mathrm{Tr}}(YY^{*})^{p}]$ may be expressed as a sum over all pairings as in Lemma 4.2 while in the corresponding expression for $\mathbf{E}[\mathop{\mathrm{Tr}}(ZZ^{*})^{p}]$ the sum is taken only over those pairings that pair even with odd indices (cf. Lemma 2.5).
The recursion follows for $p\geq 2$ by applying Corollary 5.5 to $A_{p+1}^{\prime}$ on the right-hand side of the recursion of Theorem 5.6, and a tedious but straightforward simplification of the resulting expression. That the same expression remains valid for $p=1$ can be verified directly using the explicit values for $B_{0},B_{1},B_{2},A_{0}^{\prime},A_{1}^{\prime},A_{2}^{\prime}$ that are given in Theorem 5.6 and Corollary 5.5, respectively. ∎
We will always use Lemma 5.12 for $1\leq k\leq p$ in the simplified form
$D_{k+1}\leq 2(c+1)D_{k}-\bigg{(}(c-1)^{2}-\frac{4p^{2}}{n^{2}}\bigg{)}D_{k-1}+\frac{5(k-1)^{2}}{n^{2}}A_{k-1},$ (5.4)
where we used that $4k(k-1)+1\leq 4p^{2}$, $(3k-1)(k-1)\leq 5(k-1)^{2}$, and $A_{k-1}^{\prime}\leq A_{k-1}$. As in our analysis of the complex Wishart moments, the following analysis will depend on the sign of the second term on the right-hand side.
5.3.1. The case $2p<(c-1)n$
In this subsection, we fix $1\leq p<\frac{1}{2}(c-1)n$ and let
$\mu:=c+1+2\sqrt{c+\frac{p^{2}}{n^{2}}},\qquad\qquad\bar{\mu}:=c+1-2\sqrt{c+\frac{p^{2}}{n^{2}}}.$
Note that $2(c+1)=\mu+\bar{\mu}$ and $(c-1)^{2}-\frac{4p^{2}}{n^{2}}=\mu\bar{\mu}$. Thus $\bar{\mu}>0$ as $2p<(c-1)n$.
Lemma 5.13.
We have
$D_{p}\leq\frac{\mu^{p}}{n}+\sum_{l=1}^{p-1}\frac{\mu^{p-l}-\bar{\mu}^{p-l}}{\mu-\bar{\mu}}\,\frac{5(l-1)^{2}}{n^{2}}A_{l-1}.$
Proof.
We can equivalently write (5.4) as
$D_{k+1}-\mu D_{k}\leq\bar{\mu}(D_{k}-\mu D_{k-1})+\frac{5(k-1)^{2}}{n^{2}}A_{k-1}$
for $1\leq k\leq p$. Iterating this inequality yields
$D_{k+1}-\mu D_{k}\leq\sum_{l=1}^{k}\bar{\mu}^{k-l}\,\frac{5(l-1)^{2}}{n^{2}}A_{l-1}$
for $1\leq k\leq p$, where we used that $\bar{\mu}>0$ and $D_{1}-\mu D_{0}<0$. Iterating again yields
$D_{k+1}$ $\leq\mu^{k}D_{1}+\sum_{r=1}^{k}\mu^{k-r}\sum_{l=1}^{r}\bar{\mu}^{r-l}\,\frac{5(l-1)^{2}}{n^{2}}A_{l-1}$
$=\mu^{k}D_{1}+\sum_{l=1}^{k}\frac{\mu^{k-l+1}-\bar{\mu}^{k-l+1}}{\mu-\bar{\mu}}\,\frac{5(l-1)^{2}}{n^{2}}A_{l-1}$
for $1\leq k\leq p$. The conclusion follows as $\frac{D_{1}}{\mu}\leq\frac{1}{n}$. ∎
We can now use Proposition 5.9 to estimate $D_{p}$.
Proposition 5.14.
For $1\leq p<\frac{1}{2}(c-1)n$, we have
$D_{p}\lesssim\frac{1}{n}(\sqrt{c}+1)^{2p}\bigg{(}1+\frac{p^{2}}{c^{3/2}n^{2}}\bigg{)}^{p}.$
Proof.
We can estimate as in the proof of Proposition 5.9
$\mu\leq(\sqrt{c}+1)^{2}\bigg{(}1+\frac{p^{2}}{c^{3/2}n^{2}}\bigg{)}$
Thus the only difficulty is to bound the second term on the right-hand side of the inequality of Lemma 5.13. To this end, we estimate
$\sum_{l=1}^{p-1}\frac{\mu^{p-l}-\bar{\mu}^{p-l}}{\mu-\bar{\mu}}\,\frac{5(l-1)^{2}}{n^{2}}A_{l-1}\\ \lesssim(\sqrt{c}+1)^{2(p-1)}\bigg{(}1+\frac{p^{2}}{c^{3/2}n^{2}}\bigg{)}^{p-1}c^{1/4}\sum_{r=1}^{p-2}\bigg{(}\frac{1+\frac{p^{2}}{4c^{3/2}n^{2}}}{1+\frac{p^{2}}{c^{3/2}n^{2}}}\bigg{)}^{r}\frac{r^{1/2}}{n^{2}}$
using $\mu-\bar{\mu}\geq 4\sqrt{c}$, the above inequality for $\mu$, and Proposition 5.9. We now split the sum into parts with $r\leq n^{2/3}\sqrt{c}$ and $r>n^{2/3}\sqrt{c}$. For the first part, we have
$\sum_{r=1}^{\lfloor n^{2/3}\sqrt{c}\rfloor}\bigg{(}\frac{1+\frac{p^{2}}{4c^{3/2}n^{2}}}{1+\frac{p^{2}}{c^{3/2}n^{2}}}\bigg{)}^{r}\frac{r^{1/2}}{n^{2}}\leq\frac{1}{n^{2}}\sum_{r=1}^{\lfloor n^{2/3}\sqrt{c}\rfloor}r^{1/2}\lesssim\frac{c^{3/4}}{n}.$
For the second part, we have
$\sum_{r=\lfloor n^{2/3}\sqrt{c}\rfloor+1}^{p-2}\bigg{(}\frac{1+\frac{p^{2}}{4c^{3/2}n^{2}}}{1+\frac{p^{2}}{c^{3/2}n^{2}}}\bigg{)}^{r}\frac{r^{1/2}}{n^{2}}$ $\leq\frac{c^{3/4}}{n}\sum_{r=\lfloor n^{2/3}\sqrt{c}\rfloor+1}^{p-2}\bigg{(}\frac{1+\frac{p^{2}}{4c^{3/2}n^{2}}}{1+\frac{p^{2}}{c^{3/2}n^{2}}}\bigg{)}^{r}\frac{r^{2}}{c^{3/2}n^{2}}$
$\leq\frac{c^{3/4}}{n}\frac{p^{2}}{c^{3/2}n^{2}}\sum_{r=0}^{\infty}\bigg{(}\frac{1+\frac{p^{2}}{4c^{3/2}n^{2}}}{1+\frac{p^{2}}{c^{3/2}n^{2}}}\bigg{)}^{r}$
$=\frac{4}{3}\frac{c^{3/4}}{n}\bigg{(}1+\frac{p^{2}}{c^{3/2}n^{2}}\bigg{)}.$
Combining the above estimates with Lemma 5.13 readily yields the conclusion. ∎
5.3.2. The case $2p\geq(c-1)n$
We need the following counterpart of Lemma 5.13.
Lemma 5.15.
For $p\geq 1$ such that $2p\geq(c-1)n$, we have
$D_{p}\leq C^{p-1}D_{1}+\sum_{l=1}^{p-1}C^{p-1-l}\,\frac{5(l-1)^{2}}{n^{2}}A_{l-1}$
with
$C:=(\sqrt{c}+1)^{2}\bigg{(}1+\frac{8p^{2}}{c^{2}n^{2}}\bigg{)}.$
Proof.
The assumption $2p\geq(c-1)n$ implies
$2(c+1)\leq(\sqrt{c}+1)^{2}+\frac{4p^{2}}{cn^{2}}\leq(\sqrt{c}+1)^{2}\bigg{(}1+\frac{4p^{2}}{c^{2}n^{2}}\bigg{)}$
as in the proof of Proposition 5.11. Thus (5.4) yields
$D_{k+1}\leq(\sqrt{c}+1)^{2}\bigg{(}1+\frac{4p^{2}}{c^{2}n^{2}}\bigg{)}D_{k}+(\sqrt{c}+1)^{2}\frac{4p^{2}}{cn^{2}}D_{k-1}+\frac{5(k-1)^{2}}{n^{2}}A_{k-1}$ (5.5)
for $1\leq k\leq p$. We will show by induction that for $1\leq k\leq p$
$D_{k}\leq C^{k-1}D_{1}+\sum_{l=1}^{k-1}C^{k-1-l}\,\frac{5(l-1)^{2}}{n^{2}}A_{l-1}.$
The claim is trivial for $k=1$, and follows for $k=2$ from (5.5) using $cD_{0}\leq D_{1}$. On the other hand, if the claim has been verified up to $k$, then (5.5) yields
$D_{k+1}\leq(\sqrt{c}+1)^{2}\bigg{[}\bigg{(}1+\frac{4p^{2}}{c^{2}n^{2}}\bigg{)}C+\frac{4p^{2}}{cn^{2}}\bigg{]}\bigg{[}C^{k-2}D_{1}+\sum_{l=1}^{k-2}C^{k-2-l}\,\frac{5(l-1)^{2}}{n^{2}}A_{l-1}\bigg{]}$
$\qquad\qquad\qquad+(\sqrt{c}+1)^{2}\bigg{(}1+\frac{4p^{2}}{c^{2}n^{2}}\bigg{)}\frac{5(k-2)^{2}}{n^{2}}A_{k-2}+\frac{5(k-1)^{2}}{n^{2}}A_{k-1}$
$\leq C^{2}\bigg{[}C^{k-2}D_{1}+\sum_{l=1}^{k-2}C^{k-2-l}\,\frac{5(l-1)^{2}}{n^{2}}A_{l-1}\bigg{]}+C\,\frac{5(k-2)^{2}}{n^{2}}A_{k-2}+\frac{5(k-1)^{2}}{n^{2}}A_{k-1}$
$=C^{k}D_{1}+\sum_{l=1}^{k}C^{k-l}\,\frac{5(l-1)^{2}}{n^{2}}A_{l-1},$
concluding the proof of the claim for $k+1$. The case $k=p$ yields the result. ∎
We now proceed as in the proof of Proposition 5.14.
Proposition 5.16.
For $p\geq\frac{1}{2}(c-1)n$, we have
$D_{p}\lesssim\frac{1}{n}(\sqrt{c}+1)^{2p}\bigg{(}1+\frac{8p^{2}}{c^{3/2}n^{2}}\bigg{)}^{p}.$
Proof.
The conclusion is trivial if $p=0$. For $p\geq 1$, we obtain
$D_{p}\lesssim\frac{C^{p}}{n}+(\sqrt{c}+1)^{2(p-2)}\bigg{(}1+\frac{8p^{2}}{c^{3/2}n^{2}}\bigg{)}^{p-2}c^{3/4}\sum_{r=1}^{p-2}\bigg{(}\frac{1+\frac{2p^{2}}{c^{3/2}n^{2}}}{1+\frac{8p^{2}}{c^{3/2}n^{2}}}\bigg{)}^{r}\frac{r^{1/2}}{n^{2}}$
using Lemma 5.15, $\frac{D_{1}}{C}\leq\frac{1}{n}$, $C\leq(\sqrt{c}+1)^{2}\big{(}1+\frac{8p^{2}}{c^{3/2}n^{2}}\big{)}$, and Theorem 5.1. The conclusion now follows exactly as in the proof of Proposition 5.14. ∎
The proof of Theorem 5.2 is now immediate.
Proof of Theorem 5.2.
Combining Propositions 5.14 and 5.16 yields
$B_{p}\leq D_{p}+A_{p}\lesssim(\sqrt{c}+1)^{2p}\bigg{(}1+\frac{8p^{2}}{c^{3/2}n^{2}}\bigg{)}^{p}\bigg{(}\frac{c^{3/4}}{p^{3/2}}+\frac{1}{n}\bigg{)},$
using $A_{p}^{\prime}\leq A_{p}$ and Theorem 5.1. The result follows from the definition of $B_{p}$. ∎
Acknowledgment
This work was supported in part by NSF grants DMS-1856221 and DMS-2054565.

### v2

@@
 In the special case that $b_{ij}=1$ for all $i,j$, we recover one of the most classical models of random matrix theory: then $\|X\|^{2}=\|X^{*}X\|$ is the norm of a Wishart matrix with unit covariance.
-In this case, classical methods of random matrix theory yield the exact asymptotic behavior [11]
+In this case, classical methods of random matrix theory yield the exact asymptotic behavior [Joh01]
 $\lim_{\begin{subarray}{c}n,m\to\infty\\ m=cn\end{subarray}}\mathbf{P}\big{[}\|X\|>\sqrt{n}+\sqrt{m}+\tfrac{1}{2}(\tfrac{1}{\sqrt{n}}+\tfrac{1}{\sqrt{m}})^{1/3}s\big{]}=1-F_{1}(s),$ (1.1)
@@
 From this expression, we immediately read off that in the above asymptotic regime $\|X\|=(1+o(1))(\sqrt{n}+\sqrt{m})$ with fluctuations of order $n^{-1/6}\vee m^{-1/6}$.
-It is important to note that the scale of the fluctuations is much smaller than what is predicted by general concentration of measure principles [13], which yield fluctuations of order $O(1)$ in the present setting.
-The presence of such unexpectedly small fluctuations is sometimes called the superconcentration phenomenon [6].
+It is important to note that the scale of the fluctuations is much smaller than what is predicted by general concentration of measure principles [Led01], which yield fluctuations of order $O(1)$ in the present setting.
+The presence of such unexpectedly small fluctuations is sometimes called the superconcentration phenomenon [Cha14].
 The Wishart model is amenable to explicit computations due to its simple structure and large degree of symmetry.
@@
 $\sigma_{1}^{2}:=\max_{j\leq m}\sum_{i\leq n}b_{ij}^{2},\qquad\quad\sigma_{2}^{2}:=\max_{i\leq n}\sum_{j\leq m}b_{ij}^{2},\qquad\quad\sigma_{*}^{2}:=\max_{\begin{subarray}{c}i\leq n\\ j\leq m\end{subarray}}b_{ij}^{2}.$ (1.2)
-The following is shown in [3, Corollary 3.11].
-Theorem 1.1 ([3]).
+The following is shown in [BvH16, Corollary 3.11].
+Theorem 1.1 ([BvH16]).
 Let $n\leq m$, and let $X$ be the $n\times m$ random matrix whose entries $X_{ij}\sim N(0,b_{ij}^{2})$ are independent. Then we have
@@
 In fact, (1.3) optimally captures the phase transition for the leading order behavior of the norm of random band matrices (cf.
-[3, Corollary 4.4] in the self-adjoint case).
+[BvH16, Corollary 4.4] in the self-adjoint case).
 Such examples suggest that the large scale of the fluctuations in Theorem 1.1 is a necessary feature of any bound for random matrices with an arbitrary variance pattern.
@@
 This captures precisely the fluctuations and the upper tail of the Tracy-Widom asymptotics (1.1), up to a universal constant.
-To the best of our knowledge, the nonasymptotic bound is new even in this very special case.
+To the best of our knowledge, the nonasymptotic bound is new even in this very special case (a slightly weaker bound in this setting may be found in [LR10, Theorem 2]).
 For arbitrary variance patterns with $\sigma_{1}\leq\sigma_{2}$, Theorem 1.4 yields
@@
 Theorem 1.4 controls small deviations, that is, deviations up to the order of the mean.
-In contrast, the large deviations of $\|X\|$ are controlled by its Gaussian tail behavior [17, Corollary 3.2], and we cannot expect a qualitative improvement over Theorem 1.1 in this setting.
+In contrast, the large deviations of $\|X\|$ are controlled by its Gaussian tail behavior [LT91, Corollary 3.2], and we cannot expect a qualitative improvement over Theorem 1.1 in this setting.
 Nonetheless, the basic principle behind Theorem 1.4 gives rise to a quantitative improvement: a variant of Theorem 1.1 with optimal constants.
@@
 To explain this principle, let us begin by recalling the idea behind the proof of Theorem 1.1.
-Rather than bound the norm of the nonhomogeneous model $X$ directly, the approach of [3] compares $X$ with an i.i.d.
+Rather than bound the norm of the nonhomogeneous model $X$ directly, the approach of [BvH16] compares $X$ with an i.i.d.
 matrix of smaller dimension, whose norm can be controlled by any classical method for homogeneous random matrices.
-In particular, the main technical device in [3] shows that when normalized so that $\sigma_{*}=1$, the $2p$-moment of $X$ can be bounded by the $2p$-moment of a matrix with i.i.d.
+In particular, the main technical device in [BvH16] shows that when normalized so that $\sigma_{*}=1$, the $2p$-moment of $X$ can be bounded by the $2p$-moment of a matrix with i.i.d.
 entries of dimension $\lceil\sigma_{1}^{2}+p\rceil\times\lceil\sigma_{2}^{2}+p\rceil$.
@@
 entries.
-Thus, in contrast to the comparison principle of [3], Theorem 1.7 may be viewed as an extremum principle for random matrices.
+Thus, in contrast to the comparison principle of [BvH16], Theorem 1.7 may be viewed as an extremum principle for random matrices.
 In terms of normalized moments, this principle may be expressed as follows.
@@
 With this result in hand, the proof of Theorems 1.4 and 1.6 will reduce to the proof of analogous tail bounds for Wishart matrices.
-The latter is made possible by a beautiful method, pioneered by Ledoux [14, 15, 16], for deriving Tracy-Widom type tail bounds from sharp moment estimates.
-While such estimates were obtained by Ledoux for GUE [14] and GOE [16] matrices, sharp moment bounds for Wishart matrices do not appear to be known in the literature.
+The latter is made possible by a beautiful method, pioneered by Ledoux [Led04, Led07, Led09], for deriving Tracy-Widom type tail bounds from sharp moment estimates.111 Tail bounds for Wigner and Wishart matrices have also been proved by methods that are not based on moments, such as in [Aub05, LR10].
+However, as Theorem 1.7 only provides a comparison principle for moments and not for tail probabilities, the moment approach is essential here.
+While such estimates were obtained by Ledoux for GUE [Led04] and GOE [Led09] matrices, sharp moment bounds for Wishart matrices do not appear to be known in the literature.
 Such bounds are obtained here in section 5.
 The Wishart model is considerably more delicate than the GUE and GOE models due to the fact that we require bounds that hold uniformly for all dimensions $n,m$, not just in the asymptotic regime (1.1) where $m=cn$ are proportional.
-Let us emphasize that, despite the apparent similarity between Theorem 1.7 and the comparison principles of [3], the proofs of these results are completely different.
-The method of [3] uses a classical combinatorial expression for the moments as a sum over equivalence classes of even closed walks, and estimates each term separately.
+Let us emphasize that, despite the apparent similarity between Theorem 1.7 and the comparison principles of [BvH16], the proofs of these results are completely different.
+The method of [BvH16] uses a classical combinatorial expression for the moments as a sum over equivalence classes of even closed walks, and estimates each term separately.
 In contrast, the proof of Theorem 1.7 expresses the moments of a Gaussian random matrix as a sum over pairings by the Wick formula, and then uses an iterative procedure to reduce each term to a noncrossing pairing.
@@
 The proof of the the extremum principle is much more involved in this case, and shares the same difficulties as the proof of the extremum principle for non-self-adjoint models.
-However, for both complex and real self-adjoint models, we can directly invoke the moment estimates of Ledoux [14, 15, 16] to obtain tail bounds.
+However, for both complex and real self-adjoint models, we can directly invoke the moment estimates of Ledoux [Led04, Led07, Led09] to obtain tail bounds.
   3. $\bullet$
@@
   1. 1.
-    The extremum principles of [3] and of this paper compare the spectral statistics of $X$ with those of an i.i.d. random matrix $Y$ of smaller dimension.
+    The extremum principles of [BvH16] and of this paper compare the spectral statistics of $X$ with those of an i.i.d. random matrix $Y$ of smaller dimension.
   2. 2.
-    The intrinsic freeness principle of [1] compares the spectral statistics of $X$ with those of a deterministic operator $X_{\rm free}$ that arises in free probability theory.
-The theory of [1] is in fact much more general than the setting of this paper, in that it captures both random matrices with dependent entries and more general spectral statistics.
-However, the following special case may be directly compared to the bounds of this paper: if $X$ is an $n\times m$ random matrix with independent Gaussian entries $X_{ij}\sim N(0,b_{ij}^{2})$ and $n\leq m$, then [1, Corollary 2.2] yields
+    The intrinsic freeness principle of [BBV23] compares the spectral statistics of $X$ with those of a deterministic operator $X_{\rm free}$ that arises in free probability theory.
+The theory of [BBV23] is in fact much more general than the setting of this paper, in that it captures both random matrices with dependent entries and more general spectral statistics.
+However, the following special case may be directly compared to the bounds of this paper: if $X$ is an $n\times m$ random matrix with independent Gaussian entries $X_{ij}\sim N(0,b_{ij}^{2})$ and $n\leq m$, then [BBV23, Corollary 2.2] yields
 […]
-for every $\varepsilon>0$ (this may be shown as in [1, Lemma 3.1]). As
+for every $\varepsilon>0$ (this may be shown as in [BBV23, Lemma 3.1]). As
 $\|X_{\rm free}\|\leq\sigma_{1}+\sigma_{2}$
-by [1, Lemma 2.5], the intrinsic freeness principle readily yields a slightly weaker form of (1.3).
+by [BBV23, Lemma 2.5], the intrinsic freeness principle readily yields a slightly weaker form of (1.3).
 Note that the second term in (1.6) is even larger than that of (1.3), while the results of this paper yield a much smaller second-order term.
@@
 Proof.
-We can apply [1, Lemma 3.2 and Remark 2.6] to obtain
+We can apply [BBV23, Lemma 3.2 and Remark 2.6] to obtain
 $\|X_{\rm free}\|\geq 2\sum_{i}\sqrt{w_{i}\sum_{j}b_{ij}^{2}v_{j}}+2\sum_{j}\sqrt{w_{i}\sum_{i}b_{ij}^{2}v_{j}}$
@@
 as $n\to\infty$ (cf.
-[1, 2] and [19, §2.1]).
-Moreover, it is shown in [19, Theorem 3] that when $\delta>1$, the fluctuations of $\|X\|$ are of order $\sigma_{*}\sim n^{-1/2}$.
+[BBV23, BCSV23] and [Pau07, §2.1]).
+Moreover, it is shown in [Pau07, Theorem 3] that when $\delta>1$, the fluctuations of $\|X\|$ are of order $\sigma_{*}\sim n^{-1/2}$.
 This shows that we cannot replace $\sigma_{1}+\sigma_{2}$ by $\|X_{\rm free}\|$ in Theorem 1.4, as that would imply a much smaller bound on the fluctuations of order $\sigma_{*}^{4/3}$.
 Example 1.11 illustrates in a particularly clear manner that the extremum and intrinsic freeness principles capture fundamentally different mathematical phenomena.
-This may appear somewhat surprising, given that both the proofs in this paper and those of [1] are based on the control of crossings in the Wick formula (in the latter case, this idea is due to [22]).
+This may appear somewhat surprising, given that both the proofs in this paper and those of [BBV23] are based on the control of crossings in the Wick formula (in the latter case, this idea is due to [Tro18]).
 However, these results exploit crossings in very different ways: here we show that the contribution of each crossing to the nonhomogeneous model is bounded by that in the i.i.d.
@@
 1.2.2. Superconcentration
-The concentration of measure phenomenon [13] is a powerful tool for bounding the fluctuations of random structures.
+The concentration of measure phenomenon [Led01] is a powerful tool for bounding the fluctuations of random structures.
 While general concentration inequalities often yield optimal bounds, there are also situations where the scale of the flucutations is much smaller than is predicted by general principles.
-This phenomenon was called superconcentration in [6].
+This phenomenon was called superconcentration in [Cha14].
 When applied to the random matrix models of this paper, classical concentration inequalities ensure that the scale of the fluctuations of $\|X\|$ is at most of order $\sigma_{*}$, cf.
-[1, Corollary 4.14].
+[BBV23, Corollary 4.14].
 In contrast, the fluctuation term in Theorem 1.4, of order $\sigma_{*}^{4/3}\sigma_{1}^{-1/3}$, is often much smaller than $\sigma_{*}$.
@@
 Indeed, this bound accurately captures the well known superconcentration property of the upper tail of the norm of Wishart matrices, and extends it to a much larger class of models.
-[^1] After the completion of this paper, we learned that a weaker superconcentration property was previously obtained by entirely different methods in [9, Theorem 3.4] for self-adjoint random matrices satisfying the analogue of Lemma 1.10.
+[^2] After the completion of this paper, we learned that a weaker superconcentration property was previously obtained by entirely different methods in [EK11, Theorem 3.4] for self-adjoint random matrices satisfying the analogue of Lemma 1.10.
 Theorem 3.4 below yields a much more precise tail bound in the self-adjoint setting that is essentially optimal by the extremum principle.
@@
 More generally, it is not expected that the specific setting of Lemma 1.10 is necessary for superconcentration.
-General principles explained in [6] suggest that the spectral norm should exhibit superconcentration provided there are many singular values near the maximal one (this fails in Example 1.11, where the largest singular value is isolated from the bulk).
-Furthermore, even in the setting of Lemma 1.10, our results yield an upper bound that is sharp for block-diagonal matrices, but other such models may exhibit even smaller fluctuations (e.g., [21]).
+General principles explained in [Cha14] suggest that the spectral norm should exhibit superconcentration provided there are many singular values near the maximal one (this fails in Example 1.11, where the largest singular value is isolated from the bulk).
+Furthermore, even in the setting of Lemma 1.10, our results yield an upper bound that is sharp for block-diagonal matrices, but other such models may exhibit even smaller fluctuations (e.g., [Sod10]).
 A precise understanding of when and how much superconcentration arises for nonhomogeneous random matrices remains out of reach of any known method.
@@
 The setting of Model 1.3 requires the entries of $X$ to be symmetrically distributed and have all their moments dominated by those of the Gaussian distribution.
-However, the classical Tracy-Widom asymptotics (1.1) remain valid in a much more general setting [20]: it suffices that the entries of $X$ have zero mean, unit variance, and soft control of the higher moments.
+However, the classical Tracy-Widom asymptotics (1.1) remain valid in a much more general setting [PY14]: it suffices that the entries of $X$ have zero mean, unit variance, and soft control of the higher moments.
 It is of considerable interest to understand whether nonasymptotic bounds can also be achieved in this much more general setting.
-An extension of Theorem 1.1 along these lines was proved in [12, §4.3]: it was shown there that for any random matrix $X$ whose entries are independent and have zero mean, the statement of Theorem 1.1 remains valid if we replace the parameters (1.2) by
+An extension of Theorem 1.1 along these lines was proved in [LVY18, §4.3]: it was shown there that for any random matrix $X$ whose entries are independent and have zero mean, the statement of Theorem 1.1 remains valid if we replace the parameters (1.2) by
 $\sigma_{1}^{2}\leftarrow\max_{j}\sum_{i}\mathrm{Var}(X_{ij}),\qquad\sigma_{2}^{2}\leftarrow\max_{i}\sum_{j}\mathrm{Var}(X_{ij}),\qquad\sigma_{*}^{2}\leftarrow\max_{i,j}\|X_{ij}\|_{\infty}^{2}.$
@@
 It is natural to conjecture the validity of an analogous extension of Theorem 1.4 and of the other main results of this paper.
-Such results would be of particular interest in applications, e.g., to random graphs [4].
-Unfortunately, it is not clear whether the methods of this paper can be adapted to achieve such results (nor do they follow from the very general universality principles in [5], whose error terms are far larger than the fluctuations in Theorem 1.4).
-While much less precise, the method used in [3, 12] does not use any special properties of the Gaussian distribution and is therefore readily adapted to more general models.
+Such results would be of particular interest in applications, e.g., to random graphs [BBK20].
+Unfortunately, it is not clear whether the methods of this paper can be adapted to achieve such results (nor do they follow from the very general universality principles in [BV23], whose error terms are far larger than the fluctuations in Theorem 1.4).
+While much less precise, the method used in [BvH16, LVY18] does not use any special properties of the Gaussian distribution and is therefore readily adapted to more general models.
 In contrast, the proof of the sharp extremum principle of Theorem 1.7 is based on the Wick formula, and is therefore inherently Gaussian in nature.
@@
 $\mathbf{E}[g_{i_{1}j_{1}}^{\varepsilon_{1}}\cdots g_{i_{2p}j_{2p}}^{\varepsilon_{2p}}]=\sum_{\pi\in\mathrm{P}_{2}([2p])}\prod_{\{k,l\}\in\pi}1_{i_{k}=i_{l},j_{k}=j_{l},\varepsilon_{k}\neq\varepsilon_{l}}$
-by the classical Wick formula [18, Theorem 22.3 and Remark 22.5]. ∎
-Free probability theory suggests [18, p.
+by the classical Wick formula [NS06, Theorem 22.3 and Remark 22.5]. ∎
+Free probability theory suggests [NS06, p.
 367] that the expression in Lemma 2.5 should be dominated by noncrossing pairings.
@@
 $\sum_{i\geq j,k\geq l}\sum_{\varepsilon,\delta\in\{1,*\}}\mathop{\mathrm{tr}}U_{ij}^{\varepsilon}M_{1}U_{kl}^{\delta}M_{2}U_{ij}^{!\varepsilon}M_{3}U_{kl}^{!\delta}M_{4}\leq\sigma_{*}^{4}\mathop{\mathrm{tr}}M_{3}M_{2}M_{1}M_{4},$
-with equality when $b_{ij}=1$ for all $i,j$. Here we denote !1:=*!1:=*! 1 := * and !*:=1!*:=1! * := 1.
+with equality when $b_{ij}=1$ for all $i,j$. Here we denote !1:=!1:=*! 1 := ∗ and !:=1!*:=1! ∗ := 1.
 Proof.
@@
 Remark 2.8.
-An identity of the above form for $\mathbf{E}[\mathop{\mathrm{tr}}Y^{2p}]$ is classical in random matrix theory, where it is known as the genus expansion [18, Theorem 22.12].
+An identity of the above form for $\mathbf{E}[\mathop{\mathrm{tr}}Y^{2p}]$ is classical in random matrix theory, where it is known as the genus expansion [NS06, Theorem 22.12].
 In particular, the precise combinatorial meaning of $\ell(\pi)$ can be understood in terms of the genus of the orientable surface obtained by gluing together the edges of a regular $2p$-gon corresponding to each pair of $\pi$.
@@
 2.2. Small deviations
-We can now combine Theorem 2.2 with a method of Ledoux [15, §5.2] to obtain small deviations inequalities at the Tracy-Widom scale.
+We can now combine Theorem 2.2 with a method of Ledoux [Led07, §5.2] to obtain small deviations inequalities at the Tracy-Widom scale.
 The basis for this method is an accurate estimate on the $p$th moment of $Y$ for moderately large $p$.
-The following can be read off from [15, pp.
+The following can be read off from [Led07, pp.
 210–211].
@@
 The moment estimate of Lemma 2.9 is not accurate for large $p$: indeed, this estimate yields a subexponential bound $\mathbf{E}[\mathop{\mathrm{tr}}Y^{2p}]^{1/2p}=O(p)$ as $p\to\infty$, while the large deviations of the norms of Gaussian random matrices are in fact subgaussian.
-To obtain a large deviation inequality, we will instead employ the following estimate of Haagerup and Thorbjørnsen [10, Eq.
+To obtain a large deviation inequality, we will instead employ the following estimate of Haagerup and Thorbjørnsen [HT03, Eq.
 (3.5)].
@@
 $\mathbf{E}[\mathop{\mathrm{tr}}e^{tY}]\leq e^{2\sqrt{d}\,t+t^{2}/2}.$
-This yields the following.
-Proposition 2.12.
-For $X$ as in Model 2.1 and $\varepsilon\geq 0$, we have
-$\mathbf{P}\big{[}\|X\|>2\sqrt{\sigma^{2}+\sigma_{*}^{2}}+\sigma_{*}\varepsilon\big{]}\leq 2n\,e^{-\varepsilon^{2}/2}.$
-Proof.
+We can now prove Theorem 2.4.
+Proof of Theorem 2.4.
 Suppose first that $\sigma_{*}=1$, and let $d=\lceil\sigma^{2}\rceil$.
@@
 for all $\varepsilon\geq 0$, where we used $d\leq\sigma^{2}+1$.
-This concludes the proof for $\sigma_{*}=1$.
-For general $\sigma_{*}$, it suffices to apply the above bound to the random matrix $\frac{X}{\sigma_{*}}$.
-∎
-Theorem 2.4 follows readily.
-Proof of Theorem 2.4.
-The conclusion follows immediately from Proposition 2.12 using that $\sqrt{\sigma^{2}+\sigma_{*}^{2}}\leq\sigma+\frac{\sigma_{*}^{2}}{2\sigma}\leq\sigma+\frac{\sigma_{*}}{2}$. ∎
+For general $\sigma_{*}$, applying the above bound to the random matrix $\frac{X}{\sigma_{*}}$ yields
+$\mathbf{P}\big{[}\|X\|>2\sqrt{\sigma^{2}+\sigma_{*}^{2}}+\sigma_{*}\varepsilon\big{]}\leq 2n\,e^{-\varepsilon^{2}/2}$
+for all $\varepsilon\geq 0$. To conclude, it remains to use that $\sqrt{\sigma^{2}+\sigma_{*}^{2}}\leq\sigma+\frac{\sigma_{*}^{2}}{2\sigma}\leq\sigma+\frac{\sigma_{*}}{2}$. ∎
 3. The symmetric case
@@
 whenever $b_{ii}=\sigma_{*}$ and $\xi_{ii}\sim N(0,1)$ for all $i$ (as then $X_{ii}=\sqrt{2}\sigma_{*}\xi_{ii}$).
-Moreover, it follows from [17, Corollary 3.2] that the tail bounds of both Theorems 2.4 and 3.5 cannot be improved in general for $t\to\infty$ (even in models where $b_{ii}=0$ for all $i$).
+Moreover, it follows from [LT91, Corollary 3.2] that the tail bounds of both Theorems 2.4 and 3.5 cannot be improved in general for $t\to\infty$ (even in models where $b_{ii}=0$ for all $i$).
 3.1. Extremum principle
@@
 For any $M_{1},M_{2},M_{3}$ with nonnegative entries, Lemma 3.11(a) yields
-$\sum_{i\geq j}\sum_{k\geq l}(H_{ij}M_{1}H_{kl}M_{2}H_{ij}M_{3}H_{kl})_{rs}\\ \leq\sigma_{*}^{4}\big{(}M_{3}M_{2}M_{1}+M_{3}M_{1}^{*}M_{2}^{*}+M_{2}^{*}M_{2}^{*}M_{1}+\mathrm{Tr}[M_{1}^{*}M_{3}]M_{2}^{*}\big{)}_{rs}$
+$\sum_{i\geq j}\sum_{k\geq l}(H_{ij}M_{1}H_{kl}M_{2}H_{ij}M_{3}H_{kl})_{rs}\\ \leq\sigma_{*}^{4}\big{(}M_{3}M_{2}M_{1}+M_{3}M_{1}^{*}M_{2}^{*}+M_{2}^{*}M_{3}^{*}M_{1}+\mathrm{Tr}[M_{1}^{*}M_{3}]M_{2}^{*}\big{)}_{rs}$
 with equality if $b_{ij}=1$ for all $i,j$. We can therefore write
@@
 By the type II assumption, we can decompose $\pi\backslash\{\{a,c\},\{b,d\}\}=\pi_{I}\cup\pi_{J}$ where $\pi_{I}\in\mathrm{P}_{2}(I)$, $\pi_{J}\in\mathrm{P}_{2}(J)$ with
-$I:=(a,b)\cup(c,d),\qquad\quad J:=(1,a)\cup(b,c)\cup(d,2p).$
+$I:=(a,b)\cup(c,d),\qquad\quad J:=[1,a)\cup(b,c)\cup(d,2p].$
 Denote by $\bm{i}_{I}:=(i_{s})_{s\in I}$, with the analogous notation for $\bm{j}$ and $J$. Then
@@
 The main difference is that now $Y$ is a GOE matrix rather than a GUE matrix.
-However, a suitable moment estimate for GOE matrices was also obtained by Ledoux [16, Theorem 8].
+However, a suitable moment estimate for GOE matrices was also obtained by Ledoux [Led09, Theorem 8].
 Lemma 3.15 (Ledoux).
@@
 We clearly have $\mathop{\mathrm{tr}}e^{tY}\leq e^{t\lambda_{1}(Y)}$, where $\lambda_{1}(Y)$ denotes the largest eigenvalue of $Y$.
-It is shown in [8, Theorem 2.11] that $\mathbf{E}[\lambda_{1}(Y)]\leq 2\sqrt{d}$.
-On the other hand, the Gaussian log-Sobolev inequality implies by [13, eq.
-(5.8)] and [8, §2.2] that $\lambda_{1}(Y)$ is a $\sqrt{2}$-subgaussian random variable, that is, that
+It is shown in [DS01, Theorem 2.11] that $\mathbf{E}[\lambda_{1}(Y)]\leq 2\sqrt{d}$.
+On the other hand, the Gaussian log-Sobolev inequality implies by [Led01, eq.
+(5.8)] and [DS01, §2.2] that $\lambda_{1}(Y)$ is a $\sqrt{2}$-subgaussian random variable, that is, that
 $\mathbf{E}\big{[}e^{t\{\lambda_{1}(Y)-\mathbf{E}[\lambda_{1}(Y)]\}}\big{]}\leq e^{t^{2}}$
@@
 We clearly have $\mathop{\mathrm{tr}}e^{t(YY^{*})^{1/2}}\leq e^{t\|Y\|}$.
-It is shown in [8, Theorem 2.13] that $\mathbf{E}\|Y\|\leq\sqrt{d_{1}}+\sqrt{d_{2}}$.
-The Gaussian log-Sobolev inequality implies by [13, eq.
-(5.8)] and [8, §2.2] that $\|Y\|$ is a $1$-subgaussian random variable, that is, that
+It is shown in [DS01, Theorem 2.13] that $\mathbf{E}\|Y\|\leq\sqrt{d_{1}}+\sqrt{d_{2}}$.
+The Gaussian log-Sobolev inequality implies by [Led01, eq.
+(5.8)] and [DS01, §2.2] that $\|Y\|$ is a $1$-subgaussian random variable, that is, that
 $\mathbf{E}\big{[}e^{t\{\|Y\|-\mathbf{E}\|Y\|\}}\big{]}\leq e^{t^{2}/2}$
 for all $t$. Combining these facts yields the conclusion. ∎
-This yields the following.
-Proposition 4.13.
-For $X$ as in Model 1.3 with $n\leq m$, we have
-$\mathbf{P}\big{[}\|X\|>{\textstyle\sqrt{\sigma_{1}^{2}+\sigma_{*}^{2}}+\sqrt{\sigma_{2}^{2}+\sigma_{*}^{2}}}+\sigma_{*}\varepsilon\big{]}\leq 2n\,e^{-\varepsilon^{2}/2}$
-for all $\varepsilon\geq 0$.
-Proof.
-Suppose first that $\sigma_{*}=1$, and let $d_{1}=\lceil\sigma_{1}^{2}\rceil$ and $d_{2}=\lceil\sigma_{2}^{2}\rceil$. Then
+We can now prove Theorem 1.6.
+Proof of Theorem 1.6.
+Suppose that $\sigma_{*}=1$, and let $d_{1}=\lceil\sigma_{1}^{2}\rceil$, $d_{2}=\lceil\sigma_{2}^{2}\rceil$. Then
 $\frac{1}{2}\,\mathbf{E}\big{[}\mathop{\mathrm{tr}}e^{t(XX^{*})^{1/2}}\big{]}$ $\leq\mathbf{E}[\mathop{\mathrm{tr}}\cosh(t(XX^{*})^{1/2})]$
-$\leq\mathbf{E}[\mathop{\mathrm{tr}}\cosh(t(YY^{*})^{1/2})]\leq\mathbf{E}\big{[}\mathop{\mathrm{tr}}e^{t(YY^{*})^{1/2}}\big{]},$
+$\leq\mathbf{E}[\mathop{\mathrm{tr}}\cosh(t(YY^{*})^{1/2})]\leq\mathbf{E}\big{[}\mathop{\mathrm{tr}}e^{t(YY^{*})^{1/2}}\big{]}$
 by Theorem 4.1, where we used that the Taylor expansion of the hyperbolic cosine only has terms of even degree. As $e^{t\|X\|}\leq n\mathop{\mathrm{tr}}e^{t(XX^{*})^{1/2}}$, we obtain
@@
 $\mathbf{P}\big{[}\|X\|>{\textstyle\sqrt{\sigma_{1}^{2}+1}+\sqrt{\sigma_{2}^{2}+1}}+\varepsilon\big{]}\leq 2n\,e^{-\varepsilon^{2}/2}$
-for all $\varepsilon\geq 0$. This concludes the proof for $\sigma_{*}=1$. For general $\sigma_{*}$, it suffices to apply the above bound to the random matrix $\frac{X}{\sigma_{*}}$. ∎
-Theorem 1.6 follows readily.
-Proof of Theorem 1.6.
-The conclusion follows immediately from Proposition 4.13 using that $\sqrt{\sigma_{1}^{2}+\sigma_{*}^{2}}+\sqrt{\sigma_{2}^{2}+\sigma_{*}^{2}}\leq\sigma_{1}+\sigma_{2}+\frac{\sigma_{*}^{2}}{2\sigma_{1}}+\frac{\sigma_{*}^{2}}{2\sigma_{2}}\leq\sigma_{1}+\sigma_{2}+\sigma_{*}$.
+for all $\varepsilon\geq 0$.
+For general $\sigma_{*}$, applying the above bound to the random matrix $\frac{X}{\sigma_{*}}$ yields
+$\mathbf{P}\big{[}\|X\|>{\textstyle\sqrt{\sigma_{1}^{2}+\sigma_{*}^{2}}+\sqrt{\sigma_{2}^{2}+\sigma_{*}^{2}}}+\sigma_{*}\varepsilon\big{]}\leq 2n\,e^{-\varepsilon^{2}/2}$
+for all $\varepsilon\geq 0$.
+To conclude the proof, it remains to use that $\sqrt{\sigma_{1}^{2}+\sigma_{*}^{2}}+\sqrt{\sigma_{2}^{2}+\sigma_{*}^{2}}\leq\sigma_{1}+\sigma_{2}+\frac{\sigma_{*}^{2}}{2\sigma_{1}}+\frac{\sigma_{*}^{2}}{2\sigma_{2}}\leq\sigma_{1}+\sigma_{2}+\sigma_{*}$.
 ∎
@@
 $\mathbf{E}[\mathop{\mathrm{tr}}(YY^{*})^{p}]\lesssim n^{p}(\sqrt{c}+1)^{2p}\bigg{(}1+\frac{8p^{2}}{c^{3/2}n^{2}}\bigg{)}^{p}\bigg{(}\frac{c^{3/4}}{p^{3/2}}+\frac{1}{n}\bigg{)}.$
-A moment estimate for complex Wishart matrices was previously obtained by Ledoux in [14, p.
+A moment estimate for complex Wishart matrices was previously obtained by Ledoux in [Led04, p.
 201].
 However, the constants in Ledoux’ estimate depend in an unspecified manner on the aspect ratio $c$, making it unsuitable for the purposes of this paper.
-[^3] The argument of [14, p.
+[^4] The argument of [Led04, p.
 201] also contains a further issue, that the recursion in eq.
@@
 First, the argument of section 4.2 shows that these moment estimates yield tail bounds as in Theorem 1.4, which match the exact Tracy-Widom asymptotics (1.1) with the optimal order of the fluctuations and tail behavior.
-On the other hand, if we let $n\to\infty$ with $p,c$ fixed, it is classical [18, p.
+On the other hand, if we let $n\to\infty$ with $p,c$ fixed, it is classical [NS06, p.
 368] that both $n^{-p}\mathbf{E}[\mathop{\mathrm{tr}}(YY^{*})^{p}]$ and $n^{-p}\mathbf{E}[\mathop{\mathrm{tr}}(ZZ^{*})^{p}]$ converge to the $p$-moment $\chi_{p}^{c}$ of the Marchenko–Pastur distribution.
-By using the explicit formula for its generating function [18, p.
-205], the Darboux method [23, Theorem 5.11] yields
+By using the explicit formula for its generating function [NS06, p.
+205], the Darboux method [Wil06, Theorem 5.11] yields
 $\chi_{p}^{c}=(1+o(1))\,(\sqrt{c}+1)^{2p}\,\frac{c^{1/4}(\sqrt{c}+1)}{2\sqrt{\pi}\,p^{3/2}}$
@@
 The remainder of this section is organized as follows.
-In section 5.1, we recall the recursive formulas of [10, 7] for the moments of complex and real Wishart matrices.
+In section 5.1, we recall the recursive formulas of [HT03, CMSV16] for the moments of complex and real Wishart matrices.
 Let us emphasize at the outset that the recursive formula for real Wishart moments involves complex Wishart moments, so that we must consider the latter even if one is ultimately only interested in real Wishart matrices.
@@
 The proofs of the main results of this section are based on recursive formulas for $A_{p},A_{p}^{\prime},B_{p}$ that we presently recall.
-The following recursive formula for $A_{p}$ was obtained by Haagerup and Thorbjørnsen [10, Theorem 8.2].
+The following recursive formula for $A_{p}$ was obtained by Haagerup and Thorbjørnsen [HT03, Theorem 8.2].
 Theorem 5.4 (Haagerup-Thorbjørnsen).
@@
 The reason that we consider the modified moments $A_{p}^{\prime}$ is that they appear in the following moment recursion for $B_{p}$ due to Cunden et al.
-[7, Theorem 3.5] (the values of $B_{1},B_{2}$ stated here are readily obtained by a straightforward computation).
+[CMSV16, Theorem 3.5] (the values of $B_{1},B_{2}$ stated here are readily obtained by a straightforward computation).
 Theorem 5.6 (Cunden et al.).
@@
 5.3. Real Wishart moments
-Taking inspiration from [16], we define
+Taking inspiration from [Led09], we define
 $D_{p}:=B_{p}-A_{p}^{\prime}.$
@@
 Acknowledgment
-This work was supported in part by NSF grants DMS-1856221 and DMS-2054565.
+This work was supported in part by NSF grants DMS-1856221, DMS-2054565, and DMS-2347954. We thank the referee for helpful suggestions.